### v1

# Bounding the parameter $\beta$ of a distance-regular graph with classical parameters

Abstract
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfying $b\geq 2$ and $D\geq 3$. Let $r=1+b+b^{2}+\cdots+b^{D-1}$. In 1999, K. Metsch showed that there exists a positive constant $C(\alpha,b)$ only depending on $\alpha$ and $b$, such that if $\beta\geq C(\alpha,b)r^{2}$, then either $\Gamma$ is a Grassmann graph or a bilinear forms graph.
In this work, we show that for $b\geq 2$ and $D\geq 3$, then there exists a constant $C_{1}(\alpha,b)$ only depending on $\alpha$ and $b$, such that if $\beta\geq C_{1}(\alpha,b)r$, then either $\Gamma$ is a Grassmann graph, a bilinear forms graph, or a geometric distance-regular graph whose local graphs are the $\alpha$-clique extensions of a $\frac{\beta}{\alpha}\times r$-grid.
Key Words: distance-regular graphs, classical parameters, eigenvalues, cliques
2010 Mathematics Subject Classification. 05E30E-mail addresses: lch1994@mail.ustc.edu.cn (C. Lv), koolen@ustc.edu.cn (J.H. Koolen).
1 Introduction
For undefined notions and more details see [1] and [3]. We introduce distance-regular graphs with classical parameters in the next section. In this paper, we study distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$. It is known that the parameter $b$ is an integer such that $b\neq 0,-1$ (Proposition 4). C. Weng [19] classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\leq-2$ under weak conditions, see Remark 2 below. In this paper we consider the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\geq 1$. The distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b=1$ were classified by Terwilliger, see [1, Theorem 6.1.1]. In [13], Koolen et al. showed that for $b\geq 2$ and $D$ much larger than $b$, the parameter $\alpha$ is bounded by $b(b+1)^{2}+1$. This was improved in [11], where Koolen et al. showed that for $b\geq 2$ and $D\geq 9$, the parameter $\alpha$ is bounded by $b^{2}(b+1)+1$. Moreover, they showed that for $b=2$ and $D\geq 12$, the parameter $\alpha$ is bounded by $2$. Let $r=1+b+\cdots+b^{D-1}$. In [17, Corollary 1.3], Metsch showed that if $b\geq 2$ and $\beta\geq C(\alpha,b)r^{2}$, then the graph must be a bilinear forms graph or a Grassmann graph, where $C(\alpha,b)$ is a constant only depending on $\alpha$ and $b$. Moreover, in the same paper, he showed that, if $\alpha=b-1$, then $\beta<\frac{1}{2b-1}(2b^{4}+2b^{3}+2b^{2}+b-1)r$ or the graph is a bilinear forms graph, and if $\alpha=b$, then $\beta<\frac{8}{3}(b^{2}+2b)r$ or the graph is a Grassmann graph.
In this paper we show that if $b\geq 2$ and $\beta\geq C_{1}(\alpha,b)r$, where $C_{1}(\alpha,b)$ is a constant only depending on $\alpha$ and $b$, then the graph must be a bilinear forms graph, a Grassmann graph, or $\Gamma$ is a geometric distance-regular graph whose local graphs are the $\alpha$-clique extensions of a $\frac{\beta}{\alpha}\times r$-grid, $\alpha$ is an integer satisfying $21\leq\alpha\leq b-$ 2.
Our main result is the following.
Theorem 1
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$, such that $b\geq 1$ and $D\geq 3$. Let $r=1+b+b^{2}+\cdots+b^{D-1}$. Then one of the following holds:
  1. (1)
    $\Gamma$ is a Johnson graph;
  2. (2)
    $\Gamma$ is a Hamming or a Doob graph;
  3. (3)
    $\Gamma$ is a halved cube;
  4. (4)
    $\Gamma$ is the Gosset graph with intersection array $\{27,10,1;1,10,27\}$;
  5. (5)
    $\Gamma$ is a Grassmann graph;
  6. (6)
    $\Gamma$ is a bilinear forms graph;
  7. (7)
    $b\geq 2$ and $\beta<\max\{2\frac{b+2}{2b+3}r(b+2)(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\};$
  8. (8)
    $b\geq 3$ and $\Gamma$ is a geometric distance-regular graph whose local graphs are the $\alpha$-clique extensions of a $\frac{\beta}{\alpha}\times r$-grid, where $\alpha$ is an integer satisfying $21\leq\alpha\leq b-$ 2. In this case we also have that $\frac{b(b+1)}{\alpha+1}$ is an integer.
Remark 1
  1. (1)
    For $\alpha\in\{b-1,b\}$, [17, Corollary 1.3] gives slightly better bounds.
  2. (2)
    For $\alpha=b=2$, the bound of [17, Corollary 1.3] was slightly improved in [12].
  3. (3)
    In [5], they showed that if a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfies $\alpha=b\geq 2$, $D\geq 9$ and $\beta=b(b^{D-1}+b^{D-2}\ldots+1)$, then the graph has to be the Grassmann graph $J_{b}(2D,D)$.
  4. (4)
    In [4], they showed that if a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfies $\alpha+1=b=2$, $D\geq 3$ and $\beta=b^{D}-1$, then the graph has to be the bilinear forms graph $H_{2}(D,D)$.
  5. (5)
    For $b\geq 2$, the twisted Grassmann graphs as found by Van Dam and Koolen in [2] are not geometric and have $\alpha=b$ and $\beta=b^{2}r+b$. This shows that for $\beta\leq b^{2}r+b$ we do not need to have Item (8) of Theorem 1.
Our proof closely follows the ideas of Metsch as given in [17]. We do not assume that $\alpha$ is an integer. We obtain this once we show that the distance-regular graph is geometric. The proof of Theorem 28 is based on the ideas of Huang [7]. The linear algebra idea in the proof of Proposition 13 was first employed in [15].
Outline of the paper: In the next section we give preliminaries and definitions. In Section 3, we discuss partial linear spaces. Our main result in this section is that we give sufficient conditions for a distance-regular graph with classical parameters to be the point graph of a partial linear space. This relies on a result of Metsch. For the point graph of a partial linear space, we propose the PLS$(\gamma)$ and SPLS$(c,s)$ properties and give sufficient conditions for a distance-regular graph to satisfy these properties. Also, in this section, we give some structural results for a distance-regular graph satisfying these properties. In Section 4, we give sufficient conditions for a distance-regular graph to be geometric. In Section 5 we introduce the dual Pasch axiom and we give sufficient conditions for a geometric distance-regular graph to satisfy the dual Pasch axiom. We also discuss distance-regular graphs that satisfy the dual Pasch axiom and provide some characterizations. In the last section we give the proof of the Theorem 1.
2 Definitions and preliminaries
All the graphs considered in this paper are finite, undirected and simple. The reader is referred to [1, 3] for more information. Let $\Gamma$ be a connected graph with vertex set $V(\Gamma)$. The distance $d(x,y)$ between two vertices $x,y\in V(\Gamma)$ is the length of a shortest path between $x$ and $y$ in $\Gamma$. The diameter $D=D(\Gamma)$ of $\Gamma$ is the maximum distance between any two vertices of $\Gamma$. For each $x\in V(\Gamma)$, let $\Gamma_{i}(x)$ be the set of vertices in $\Gamma$ at distance $i$ from $x$ ($0\leq i\leq D$). In addition, define $\Gamma_{-1}(x)=\Gamma_{D+1}(x)=\emptyset$. For the sake of simplicity, we denote $\Gamma_{1}(x)$ by $\Gamma(x)$. For any vertex $x$ of $\Gamma$, the subgraph induced on $\Gamma(x)$ is called the local graph of $\Gamma$ at $x$, and we denote it by $\Delta_{x}$. For a vertex $x$ of $\Gamma$, the cardinality $|\Gamma(x)|$ of $\Gamma(x)$ is called the valency of $x$ in $\Gamma$. In particular, $\Gamma$ is regular with valency $k$ if $k=|\Gamma(x)|$ holds for all $x\in V(\Gamma)$. We also define for vertices $x,y$ of $\Gamma$ at distance $i$ $(i=1,2,\ldots,D)$ the set $C_{i}(x,y)$ by $C_{i}(x,y):=\Gamma_{i-1}(x)\cap\Gamma(y)$. For vertices $u,v$ of $\Gamma$ at distance $j$ $(j=0,1,\ldots,D-1)$, we define the set $B_{j}(u,v)$ by $B_{j}(u,v):=\Gamma_{j+1}(u)\cap\Gamma(v)$.
The adjacency matrix $A=A(\Gamma)$ of $\Gamma$ is the matrix whose rows and columns are indexed by vertices of $\Gamma$ and the ($x,y$)-entry is $1$ whenever $x$ and $y$ are adjacent and 00 otherwise. The eigenvalues of $\Gamma$ are the eigenvalues of its adjacency matrix $A$.
A clique of a graph $\Gamma$ is a set of mutually adjacent vertices of $\Gamma$. We sometimes also refer to a complete subgraph of $\Gamma$ as a clique. Let $K_{m,n}$ be the complete bipartite graph with $m$ and $n$ be the number of vertices in each partition.
A connected graph $\Gamma$ is called edge-regular with parameters $(n,k,a)$ if $\Gamma$ has $n$ vertices, is $k$-regular and any two adjacent vertices have exactly $a$ common neighbours. A connected graph $\Gamma$ is called amply-regular with parameters $(n,k,a,c)$ if $\Gamma$ is edge-regular with parameters $(n,k,a)$ and any two vertices at distance 2 have exactly $c$ common neighbours.
The $m\times n$-grid is the Cartesian product $K_{m}\Box K_{n}$ where $K_{n}$ is the complete graph on $n$ vertices.
For a positive integer $s$ and a graph $\Gamma$, the $s$-clique extension of $\Gamma$, $\tilde{\Gamma}$, is the graph obtained from $\Gamma$ by replacing each vertex $x$ with a complete graph $\tilde{X}$ having $s$ vertices, such that $u\in\tilde{X}$ is adjacent to $v\in\tilde{Y}$ $(\tilde{X}\neq\tilde{Y})$ in $\tilde{\Gamma}$ if and only if $x\sim y$ in $\Gamma$.
Let $v\geq k\geq t>0$ be integers. Let $\mathcal{P}$ be a set of $v$ elements and $\mathcal{B}$ a subset of (${\mathcal{P}}\choose k$. The pair $X=({\mathcal{P}},{\mathcal{B}})$ is called a $t$-$(v,k,\lambda)$-design if each $T\in{{\mathcal{P}}\choose t}$ lies in exactly $\lambda$ elements of $\mathcal{B}$. The sets $\mathcal{P}$ and $\mathcal{B}$ are called the point set of $X$ and the block set of $X$, respectively.
2.1 Distance-regular graphs
A graph $\Gamma$ is called distance-regular if there exist integers $b_{i},c_{i}$ $(0\leq i\leq D)$ such that for any two vertices $x,y\in V(\Gamma)$ with $d(x,y)=i$, there are precisely $c_{i}$ neighbors of $y$ in $\Gamma_{i-1}(x)$ and $b_{i}$ neighbors of $y$ in $\Gamma_{i+1}(x)$, where we define $b_{D}=c_{0}=0$. Note that in this case $C_{i}(x,y)$ contains exactly $c_{i}$ vertices for $i=1,2,\ldots,D$ and $B_{i}(x,y)$ contains exactly $b_{i}$ vertices for $i=0,1,\ldots,D-1$. In particular, any distance-regular graph is regular with valency $k:=b_{0}$. We define $a_{i}:=k-b_{i}-c_{i}$ for notational convenience. Note that $a_{i}=|\Gamma(y)\cap\Gamma_{i}(x)|$ holds for any two vertices $x,y$ with $d(x,y)=i$ $(0\leq i\leq D)$ and that the numbers $a_{i},b_{i}$ and $c_{i}$ $(0\leq i\leq D)$ are called the intersection numbers of $\Gamma$. For an eigenvalue $\theta$ of $\Gamma$, the sequence $(u_{i})_{i=0,1,...,D}$ = $(u_{i}(\theta))_{i=0,1,...,D}$ satisfying $u_{0}$ = $u_{0}(\theta)$ = $1$, $u_{1}$ = $u_{1}(\theta)$ = $\theta/k$, and
$c_{i}u_{i-1}+a_{i}u_{i}+b_{i}u_{i+1}=\theta u_{i}~{}(i=1,2,\ldots,D-1)$ (1)
is called the standard sequence corresponding to the eigenvalue $\theta$ (see [1, p.128]). For the convenience of the reader we give the proof of the following well-known lemma.
Lemma 2 (cf.[1, Proposition 4.4.1])
Let $\Gamma$ be a distance-regular graph, $\theta\neq k$ an eigenvalue of $\Gamma$, say with multiplicity $m$ and let $(u_{i})_{i}$ be the standard sequence corresponding to $\theta$. Then there exists a map $f:V(\Gamma)\rightarrow\mathbf{R}^{m}$, $x\mapsto\overline{x}$ such that the inner product $(\overline{x},\overline{y})=u_{d(x,y)}$.
The map $f$ is called the standard representation of $\Gamma$ with respect to $\theta$.
The following lemma is called the Delsarte bound. We give a proof for it, as we need some consequences later in this paper. For a set $T$ of vertices of a connected graph and a vertex $x$, we define the distance $d(x,T)$ by $d(x,T)=\min\{d(x,y)\mid y\in T\}$.
Lemma 3 (cf.[1, Proposition 4.4.6])
Let $\Gamma$ be a distance-regular graph with diameter $D\geq 2$, valency $k$ and smallest eigenvalue $\theta_{\min}$, say with multiplicity $m$. Let $(u_{i})_{i}$ be the standard sequence corresponding to $\theta_{\min}$. Then the order $c$ of a clique $C$ in $\Gamma$ is bounded by $c\leq 1+\frac{k}{-\theta_{\min}}$.
Assume that $c=1+\frac{k}{-\theta_{\min}}$. Then any vertex of $\Gamma$ has distance at most $D-1$ to $C$. Let $x$ be a vertex with $d(x,C)=j$ for some $11\leq j\leq D-$ 1. Then the number of vertices $\phi_{j}(x)$ in $C$ at distance $j$ from $x$ satisfies:
$(u_{j+1}-u_{j})\phi_{j}(x)=(1+\frac{k}{-\theta_{\min}})u_{j+1},$ (2)
and hence $\phi_{j}(x)$ does not depend on $x$, only on $j$.
Proof. Let $f:V(\Gamma)\rightarrow\mathbf{R}^{m},x\mapsto\overline{x}$ be the standard representation of $\Gamma$ with respect to $\theta_{\min}$. Let $\overline{C}=\sum\overline{u}$ where the sum is taken over all vertices $u$ of $C$. Then the Gram matrix $Gr$ with respect to $\{\overline{u}\mid u$ is a vertex of $C\}$ is the matrix that satisfies
$Gr=u_{1}J+(1-u_{1})I.$
As $u_{1}=\frac{\theta_{\min}}{k}$, and $Gr$ has row sum $1+(c-1)u_{1}$ which is non-negative, as $Gr$ is positive semidefinite, we obtain $c\leq 1+\frac{k}{-\theta_{\min}}$, and equality holds if and only if $\overline{C}=\mathbf{0}$. Assume that equality holds. Let $x$ be a vertex of $\Gamma$ such that $d(x,C)=j$. Then the inner product $(\overline{x},\overline{C})$ satisfies
$0=(\overline{x},\overline{C})=u_{j}\phi_{j}(x)+u_{j+1}(c-\phi_{j}(x)),\mbox{ for }1\leq j\leq D-1,\mbox{ and }$
$0=(\overline{x},\overline{C})=u_{D}\phi_{D}(x),\mbox{ if }j=D.$
As $u_{D}\neq 0$, we have $\phi_{D}(x)=0$. This implies that there is no vertex of $\Gamma$ has distance $D$ to $C$. This shows the lemma.
A clique of $\Gamma$ with $1+\frac{k}{-\theta_{\min}}$ vertices is called a Delsarte clique.
A distance-regular graph $\Gamma$ is called geometric if there exists a set of Delsarte cliques $\mathcal{C}$ in $\Gamma$ such that each edge lies in a unique $C\in{\mathcal{C}}$. Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$. Let $C$ a Delsarte clique and $x$ be a vertex at distance $j$ from $C$ for $10\leq j\leq D-$ 1. Define $\phi_{j}$ be the number of vertices in $C$ at distance $j$ from $x$. For two vertices $x$ and $y$ at distance $j$ for $1\leq j\leq D,$ let $\tau_{j}$ be the number of cliques in $\mathcal{C}$ containing $x$ at distance $j-1$ from $y$. The number $\tau_{j}$ does not depend on the pair $x,y$, only at their distance.
We can express the intersection numbers of $\Gamma$ in terms of the $\phi_{j}$’s and $\tau_{j}$’s as follows (see [10, Lemma 4.1]):
$c_{i}=\tau_{i}\phi_{i-1}\ \ (i=1,2,\ldots,D),\ \ b_{i}=-(\theta_{\min}+\tau_{i})(1+\frac{k}{-\theta_{\min}}-\phi_{i})\ \ (i=1,2,\ldots,D-1).$ (3)
2.2 Distance-regular graphs with classical parameters
A distance-regular graph $\Gamma$ is said to have classical parameters $(D,b,\alpha,\beta)$ if the diameter of $\Gamma$ is $D$ and the intersection numbers of $\Gamma$ can be expressed as follows:
$b_{i}=([D]-[i])(\beta-\alpha[i]),~{}~{}0\leq i\leq D-1,$
$c_{i}=[i](1+\alpha[i-1]),~{}~{}1\leq i\leq D,$
where
$[j]=\begin{cases}\frac{b^{j}-1}{b-1}&\text{if }b\neq 1,\\ {j\choose 1}&\text{if }b=1.\\ \end{cases}$
From [1, Corollary 8.4.2], we know that the eigenvalues of $\Gamma$ are
$[D-i](\beta-\alpha[i])-[i]=\frac{b_{i}}{b^{i}}-[i],~{}~{}0\leq i\leq D.$
We note that $c_{2}=(b+1)(\alpha+1)$ and that if $b\geq 1$, then the eigenvalues $\theta_{i}=\frac{b_{i}}{b^{i}}-\left[\begin{array}[]{c}i\\ 1\\ \end{array}\right](0\leq i\leq D)$ of $\Gamma$ are in the natural ordering, i.e., $k=\theta_{0}>\theta_{1}>\cdots>\theta_{D}$.
From the following result we know that the parameter $b$ of a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ is an integer.
Proposition 4 (cf.[1, Proposition 6.2.1])
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ and diameter $D\geq 3$. Then $b$ is an integer such that $b\neq 0,-1$.
The following lemma is an easy consequence of Proposition 4.
Lemma 5
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ and diameter $D\geq 3$, such that $b\geq 1$. Then $\alpha\geq 0$.
Proof. As $b$ is a positive integer and $c_{2}=(\alpha+1)(b+1)$ is clearly a positive integer, we have $\alpha(b+1)$ is an integer. Now the intersection number $c_{3}$ satisfies $c_{3}=(1+b+b^{2})(1+(1+b)\alpha)\geq 1$. This means $(1+b)\alpha+1>0$, and as it is an integer, we see $(1+b)\alpha+1\geq 1$. This shows the lemma.
The members of many of the infinite families of the known distance-regular graphs with unbounded diameter have classical parameters, like the Johnson graphs, Hamming graphs, Grassmann graphs and the bilinear forms graphs where the parameter $b$ is positive.
Remark 2
  1. (1)
    For $b\leq-2$, C. Weng [19] classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\leq-2$. His result states: If $D\geq 4$, $a_{1}\neq 0$ and $c_{2}>1$, then one of the following holds: $(1)$ $\Gamma$ is the dual polar graph ${}^{2}A_{2D-1}(-b)$, $(2)$ $\Gamma$ is the Hermitian forms graph $Her_{-b}(D)$, $(3)$ $\alpha=\frac{b-1}{2}$, $\beta=-\frac{1+b^{D}}{2}$ and $-b$ is a power of an odd prime.
  2. (2)
    Tian et al. [18] give the current state of art on distance-regular graphs with classical parameters with $b\leq-2$.
Before we look at the Delsarte bound for distance-regular graphs with classical parameters, we will determine the standard sequence for the smallest eigenvalue of such a distance-regular graph. Let $b\geq 2$ be an integer and let $[i]=\frac{b^{i}-1}{b-1}$ and $r=[D]$, as before.
Lemma 6
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $b\geq 2$ and $D\geq 2$. The standard sequence $(u_{i})_{i}$ corresponding to the smallest eigenvalue $-r$ of $\Gamma$ satisfies: $u_{0}=1$ and $u_{i}=\frac{(-1)^{i}}{\beta}\frac{(1+\alpha[1])(1+\alpha[2])\cdots(1+\alpha[i-1])}{(\beta-\alpha[1])(\beta-\alpha[2])\cdots(\beta-\alpha[i-1])}$ for $i\geq 1$.
Proof. We have $c_{i}=[i](1+\alpha[i-1])$ for $1\leq i\leq D$, $a_{i}=[i](\beta-1+\alpha(r-[i]-[i-1]))$ for $1\leq i\leq D$ and $b_{i}=(r-[i])(\beta-\alpha[i])$ for $10\leq i\leq D-$ 1. Note that $k=b_{0}=\beta r$. The $u_{i}$’s satisfy $u_{0}=1$, $u_{1}=\frac{-r}{k}=\frac{-1}{\beta}$ and $c_{i}u_{i-1}+a_{i}u_{i}+b_{i}u_{i+1}=-ru_{i}$ for $11\leq i\leq D-$ 1. The lemma follows by an easy induction.
The following result is a direct consequence of Lemmas 3 and 6 for a distance-regular graph with classical parameters.
Lemma 7
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $D\geq 3$ and $b\geq 2$. Then the order $c$ of a clique $C$ in $\Gamma$ is bounded by $c\leq\beta+1$. If equality holds, then the number $\phi_{j}$ for a vertex at distance $j$ from $C$ satisfies $\phi_{j}=1+\alpha[j]$ for $10\leq j\leq D-$ 1. In particular $\phi_{1}=1+\alpha$.
Proof. As the smallest eigenvalue $\theta_{\min}$ of $\Gamma$ is equal to $-r$ where $r=[D]$, and $k=\beta r$ we obtain that any clique $C$ with order $c$ satisfies $c\leq 1+\frac{k}{r}=1+\beta$. If equality holds then by Equation 2 we obtain $\phi_{j}(x)=1+\alpha[j]$, as the standard sequence corresponding to eigenvalue $-r$ satisfies $u_{i}=\frac{(-1)^{i}}{\beta}\frac{(1+\alpha[1])(1+\alpha[2])\cdots(1+\alpha[i-1])}{(\beta-\alpha[1])(\beta-\alpha[2])\cdots(\beta-\alpha[i-1])}$ for $i\geq 1$, by Lemma 6. This shows the lemma.
2.3 Partial linear spaces
An incidence structure is a tuple $(\mathcal{P},\mathcal{L},\mathcal{I})$ where $\mathcal{P}$ and $\mathcal{L}$ are non-empty disjoint sets and $\mathcal{I}\subseteq\mathcal{P}\times\mathcal{L}$. The elements of $\mathcal{P}$ and $\mathcal{L}$ are called points and lines, respectively. If $(p,\ell)\in\mathcal{I}$, then we say that $p$ is incident with $\ell$, or that $p$ is on the line $\ell$. The order of a point is the number of lines it is incident with and similarly for lines. The point-line incidence matrix of $(\mathcal{P},\mathcal{L},\mathcal{I})$ is the $|\mathcal{P}|\times|\mathcal{L}|$-matrix such that the $(p,\ell)$-entry is 1 if $p$ is incident with $\ell$ and 0 otherwise. If it is clear from the context what the incidence relation is, then we omit it.
A partial linear space is an incidence structure such that each pair of distinct points are both incident with at most one line.
Let $X=(\mathcal{P},\mathcal{L},\mathcal{I})$ be a partial linear space. For a non-incident pair $(x,\ell)$ of a point $x$ and line $\ell$, let $\tau(x,\ell)$ be the number of points on $\ell$ that are collinear with $x$. We define $\tau(X)$ for a partial linear space $X=(\mathcal{P},\mathcal{L},\mathcal{I})$, as the number $\tau(X)=\max\{\tau(x,\ell)\mid x\in\mathcal{P},\ell\in\mathcal{L},\mbox{ and }(x,\ell)\notin\mathcal{I}\}$.
The point graph $\Gamma$ of an incidence structure $(\mathcal{P},\mathcal{L},\mathcal{I})$ is the graph with vertex set $\mathcal{P}$ and two distinct points are adjacent if and only if they are on a common line. Note that lines are cliques in $\Gamma$. If the point graph is connected, then for a vertex $x$ and a line $\ell$, we define the distance $d(x,\ell)$ by $d(x,\ell)=\min\{d(x,y)\mid y\in\ell\}$. For two vertices $x,y$ at distance 2 in $\Gamma$, we denote by $[x,y]$ the set of lines $\ell$ through $x$ such that $d(y,u)\leq 2$ for all $u\in\ell$.
First we give the so-called claw bound. It was first proven by Metsch [16, Lemma 1.1] and was rediscovered by Godsil [6, Lemma 2.3] and Koolen and Park [14, Lemma 2].
Lemma 8
Let $\Gamma$ be an amply-regular graph with parameters $(n,k,\lambda,\mu)$. Assume that there exists a positive integer $s$ such that the following holds:
$(s+1)(\lambda+1)-k>(\mu-1){s+1\choose 2}.$
Then $\Gamma$ does not contain an induced $K_{1,s+1}$.
Now we recall the following result of Metsch. Note that the first condition is just the claw bound.
Theorem 9 (cf.[17, Result 2.1])
Let $\Gamma$ be an amply-regular graph with parameters $(n,k,\lambda,\mu)$. Assume that there exists a positive integer $s$ such that the following two conditions are satisfied:
  1. (1)
    $(s+1)(\lambda+1)-k>(\mu-1){s+1\choose 2};$
  2. (2)
    $\lambda+1>(\mu-1)(2s-1)$.
Define a line as a maximal clique with at least $\lambda+2-(\mu-1)(s-1)$ vertices. Then $X=(V(\Gamma),{\mathcal{L}},\in)$ is a partial linear space, where $\mathcal{L}$ is the set of all lines, and $\Gamma$ is the point graph of $X$. Moreover, each vertex lies on at most $s$ lines.
3 The SPLS$(s)$ Property
We say that a distance-regular graph $\Gamma$ has the PLS$(\gamma)$ property for some $\gamma\geq 3$ if $\Gamma$ is the point graph of a partial linear space $X=(V(\Gamma),{\mathcal{L}},\in)$, where the set of lines $\mathcal{L}$ are the maximal cliques with at least $\gamma$ vertices.
We have $\gamma\leq 1+\frac{k}{-\theta_{\min}}$ where $\theta_{\min}$ is the smallest eigenvalue of $\Gamma$, and equality if and only if $\Gamma$ is geometric.
Let $\Gamma$ be a distance-regular graph that is the point graph of a partial linear space $X=(V(\Gamma),{\mathcal{L}},\in)$. A vertex $x$ of $\Gamma$ is called Delsarte vertex if all the lines in ${\mathcal{L}}$ through $x$ are Delsarte cliques.
Lemma 10
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Assume that $\Gamma$ satisfies the PLS$(\gamma)$ property for some $\gamma\geq 3$. Let $r=[D]$. Then a vertex $x$ lies on at least $r$ lines and it is on exactly $r$ lines if and only if it is a Delsarte vertex.
Proof. By Lemma 7, the order $c$ of a clique $C$ in $\Gamma$ satisfies $c\leq 1+\beta$ with equality if and only if $C$ is a Delsarte clique. Now $k=\beta r$. By the definition of a partial linear space the lemma follows.
Now we prepare a lemma that will bound the parameter $\alpha$ of a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$.
Lemma 11
Let $\Gamma$ be a distance-regular graph with classical prameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Assume that $\Gamma$ satisfies the PLS$(\gamma)$ property for some $\gamma\geq 3$. Assume that $\Gamma$ contains two distinct Delsarte vertices at distance at most two. Then $\alpha\leq b$ and $\alpha$ is a non-negative integer.
Proof. Let $x,z$ be two vertices at distance two from each other such that $x$ is a Delsarte vertex and there exists a Delsarte clique $C’$ ’ containing $z$ such that $x$ has at least one neighbour in $C’$ ’ and hence $x$ has exactly $\alpha+1$ neighbours in $C’$ ’. This means that $\alpha+1$ is a positive integer. Let $r=[D]$. Let $C_{1},C_{2},\ldots,C_{r}$ be the lines through $x$, all of which are Delsarte cliques.
The vertices $x$ and $z$ have $c_{2}=(1+\alpha)(1+b)$ common neighbours and $z$ has exactly $\alpha+1$ or 0 neighbours in each of the $C_{i}$ for $i=1,2,\ldots,r$. This means that there are exactly $b+1$ $C_{i}$’s such that $z$ has $\alpha+1$ neighbours in $C_{i}$. Without loss of generality we may assume that $z$ has $\alpha+1$ neighbours in $C_{i}$ for $i=1,2,\ldots,b+1$. On the other hand $x$ has exactly $\alpha+1$ neighbours in $C’$ ’. Now the clique $C_{i}$ and $C’$ ’ intersect in at most one vertex for $i=1,2,\ldots,b+1$ and they do not intersect if $i\geq b+2$. This means that $\alpha+1\leq b+1$, and hence $\alpha\leq b$.
Let $y$ be a Delsarte vertex in $\Gamma$ at distance at most two from $x$ and distinct from $x$.
Assume first that $x\sim y$. There exists a Delsarte clique $C’$ ’ that contains $y$ but not $x$. As $C’$ ’ is a maximal clique there exists a vertex $z$ in $C’$ ’ not adjacent to $x$. So then we obtain $\alpha\leq b$ by the above.
Now assume $d(x,y)=2$. Then we take $z=y$, and we are done again.
Now we will strengthen the PLS$(\gamma)$ property as follows:
We say that a distance-regular graph $\Gamma$ has the SPLS$(c,s)$ property for some integers $c\geq 1$ and $s\geq 2$, if $\Gamma$ is the point graph of a partial linear space $X=(V(\Gamma),{\mathcal{L}},\in)$, where the set of lines $\mathcal{L}$ are the maximal cliques with at least $a_{1}+2-(c-1)(s-1)$ vertices, $a_{1}\geq(2s-1)(c-1)$, each vertex lies on at most $s$ lines and $\tau(X)\leq c$.
If $\Gamma$ has the SPLS$(c,s)$ property for some integers $c\geq 1$ and $s\geq 2$, then each line has at least $a_{1}+2-(c-1)(s-1)\geq s(c-1)+2$ vertices, and clearly $c\leq c_{2}$ holds. If $c=c_{2}$, then we write SPLS$(s)$ instead of SPLS$(c_{2},s)$.
[…]
Now we give some bounds for $s$ for a distance-regular graph satisfying the SPLS$(s)$ property.
Lemma 12
Let $\Gamma$ be a distance-regular graph with valency $k$ and smallest eigenvalue $\theta_{\min}$. If $\Gamma$ has the SPLS$(s)$ property, then $-\theta_{\min}\leq s$, and if equality holds, then $\Gamma$ is a geometric distance-regular graph. Let $\sigma$ be the smallest integer such that $\Gamma$ has the SPLS$(\sigma)$ property. Then $\sigma(a_{1}+1)-k\leq(c_{2}-1){\sigma\choose 2}$, and $\sigma\geq-\theta_{\min}$.
Proof. As each clique has at most $1+\frac{k}{-\theta_{\min}}$ vertices, it is clear that $-\theta_{\min}\leq s$ and if equality holds, then each line is a Delsarte clique and hence the graph is geometric.
Let $x$ be a vertex that lies on exactly $\sigma$ lines. Then it is easy to show that $x$ lies in an induced $K_{1,\sigma}$, by an easy induction argument as any line has at least $(c_{2}-1)\sigma+2$ vertices and that $\Gamma$ does not contain an induced $K_{1,\sigma+1}$. This shows the lemma, by applying Lemma 8.
Now we give an extra condition on distance-regular graphs with classical parameters that have the SPLS$(s)$ property.
Proposition 13
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $s$ be a positive integer and assume $\Gamma$ has the SPLS$(s)$ property. Let $r=[D]$. Then the following hold:
  1. (1)
    $s\geq r$;
  2. (2)
    more than $\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$ vertices are Delsarte vertices;
  3. (3)
    $\alpha$ is an integer such that $0\leq\alpha\leq b$ and $\beta$ is a positive integer;
  4. (4)
    There exists two vertices $x,y$ at distance 2 such that $|[x,y]|\geq b+1$.
Proof. The smallest eigenvalue of $\Gamma$ is equal to $-r$, and therefore $s\geq r$. By the SPLS$(s)$ property, let $X=(V(\Gamma),\mathcal{C},\in)$ be the partial linear space with $\Gamma$ as its point graph, where $\mathcal{C}$ is the set of maximal cliques that have at least $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2$ vertices. Note that all the Delsarte cliques of $\Gamma$ are in $\mathcal{C}$, as all cliques in $\Gamma$ have at most $\beta+1$ vertices and clearly $a_{1}+2-(s-1)(c_{2}-1)\leq\beta+1$, as otherwise $\mathcal{C}$ would be empty. Let $N$ be the point-line incidence matrix of $X$. Let $A$ be the adjacency matrix of $\Gamma$. Now $NN^{T}=A+Q$ where $Q$ is a diagonal matrix. Note that the rank of $N$ equals the rank of $A+Q$.
By Lemma 10 we see that a vertex $x$ of $\Gamma$ lies on exactly $r$ lines if and only if $x$ is a Delsarte vertex. This means that for all vertices $x$ of $\Gamma$ the entry $Q_{xx}\geq r$ and $Q_{xx}=r$ if and only if $x$ is a Delsarte vertex.
Let $\delta$ be the dimension of the null-space of $A+Q$ and $\kappa$ be the number of Delsarte vertices. Let $\mathbf{u}\neq\mathbf{0}$ such that $(A+Q)\mathbf{u}=\mathbf{0}$. We find $\mathbf{u}^{T}(A+rI)\mathbf{u}\geq 0$, as $A$ has smallest eigenvalue $-r$, and $\mathbf{u}^{T}(Q-rI)\mathbf{u}\geq\mathbf{0}$, as $Q_{xx}\geq r$ for all vertices $x$ of $\Gamma$. This implies that $A\mathbf{u}=-r\mathbf{u}$ and $Q\mathbf{u}=r\mathbf{u}$. Hence we obtain that the multiplicity $m$ of $r$ as an eigenvalue of $Q$ is at least $\delta$. Further we have that $m$ is exactly the number of vertices $x$ such that $Q_{xx}=r$. This shows that $\delta\leq\kappa$.
Now we determine the rank of $N$. This is at most the number of lines. So we need to estimate the number $\eta$ of lines. We do this by double counting the pairs $(x,\ell)$ where $x$ is a vertex of $\Gamma$ and $\ell$ is a line. Let $w:=|\{(x,\ell)\mid x\in V(\Gamma),\ell\in\mathcal{C},x\in\ell\}|.$ On the one hand we have that $w\leq s|V(\Gamma)|$ as every vertex lies on at most $s$ lines. On the other hand we find $w\geq\eta(a_{1}+2-(s-1)(c_{2}-1))$ as a line contains at least $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2$ vertices. Hence we obtain that
$\eta\leq\frac{s|V(\Gamma)|}{a_{1}+2-(s-1)(c_{2}-1)}.$
As $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2>s(c_{2}-1),$ we find that $\eta<\frac{1}{c_{2}-1}|V(\Gamma)|.$ So the rank of $N$ is less than $\frac{1}{c_{2}-1}|V(\Gamma)|$. This means that the dimension $\delta$ of the null-space of $A+Q$ is more than $|V(\Gamma)|-\frac{1}{c_{2}-1}|V(\Gamma)|=\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$. As $b\geq 2$ and $\alpha\geq 0$, by Lemma 5, we have $c_{2}=(\alpha+1)(b+1)\geq 3$. Then, as the number $\kappa$ of Delsarte vertices is at least the dimension $\delta$ and this number is in turn more than $|V(\Gamma)|-\frac{1}{c_{2}-1}|V(\Gamma)|=\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$, we find more than half the vertices are Delsarte vertices. It means that there are adjacent vertices which are both Delsarte vertices. By Lemma 11 we find that $\alpha\geq 0$ is an integer and $\alpha\leq b$. If we take $x$ as a Delsarte vertex and $y$ a vertex at distance 2 of $x$, then $y$ has $\alpha+1$ or 0 neighbours on any line through $x$. As $c_{2}=(\alpha+1)(b+1)$, we see that there are at least $b+1$ lines in $[x,y]$. This shows the proposition.
Now we give a sufficient condition for a distance-regular graph with classical parameters to have the SPLS$(s)$ property.
Theorem 14
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. If $\beta\geq\frac{1}{3}(8\alpha b+8b+5\alpha)r$, then $\Gamma$ has the SPLS$(\lfloor\frac{4r}{3}\rfloor)$ property. In particular, in this case, $\alpha$ is an integer such that $0\leq\alpha\leq b$.
Proof. Assume that $\beta\geq\frac{1}{3}(8\alpha b+8b+5\alpha)r$. We need to show that $\Gamma$ has the SPLS$(\lfloor\frac{4r}{3}\rfloor)$ property. By the definition of the SPLS$(s)$ property, we only need to show that the conditions in Theorem 9 are all satisfied for $s=\lfloor\frac{4r}{3}\rfloor$. As $\lambda+1=a_{1}+1=\beta+\alpha(r-1)$, $\mu-1=c_{2}-1=\alpha b+\alpha+b$ and $k=\beta r$, the conditions in Theorem 9 can be written as
$\beta>\frac{s+1}{2(s+1-r)}\left((\alpha b+\alpha+b)s-2\alpha(r-1)\right),$ (4)
and
$\beta>(\alpha b+b+\alpha)2s-\alpha r-(\alpha+1)b.$ (5)
Let $s_{0}:=\frac{4r}{3}$ be a rational number. Then
$\frac{s+1}{s+1-r}<\frac{s_{0}}{s_{0}-r}=4.$
So to show that Conditions (4) and (5) hold, it suffices to show that the following two conditions hold:
$\beta>2(\alpha b+b+\alpha)s_{0}-4\alpha(r-1)=\frac{4}{3}(2(\alpha b+b+\alpha)r-3\alpha(r-1))$ (6)
and
$\beta>2(\alpha b+b+\alpha)s_{0}-\alpha r-(\alpha+1)b=\frac{1}{3}(8(\alpha b+b+\alpha)r-3\alpha r-3(\alpha+1)b).$ (7)
When $\beta\geq\frac{1}{3}(8\alpha b+8b+5\alpha)r$, Conditions (6) and (7) clearly hold. This shows the theorem.
4 Geometric distance-regular graphs
In this section we will give a necessary condition for a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ which has the SPLS$(s)$ property to be geometric.
Our main result in this section is the following and we will give the proof using several claims.
Theorem 15
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $s$ be a positive integer and assume $\Gamma$ has the SPLS$(s)$ property. Let $r=[D]$. Assume that $\beta$ satisfies the following bounds:
  1. (1)
    $\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha$;
  2. (2)
    $\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$;
  3. (3)
    $\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b$.
Then $\Gamma$ is geometric.
Proof. Note that by Proposition 13, we have $s\geq r$ and $0\leq\alpha\leq b$ where $\alpha$ is an integer. Note that by (1), we have $\beta>(b^{2}+b+1)(b+1)$ as $0\leq\alpha\leq b$ and $D\geq 3$.
Now we will look at the line set $[x,y]$ for two vertices $x,y$ at distance 2. First, we show that this set has cardinality at most $12b+$ 1, and later, we will improve it to $b+1$.
Claim 1 If $\alpha\geq 1$, then$|[x,y]|<2(b+1)$ for any two vertices $x$ and $y$ of $\Gamma$ at distance 2. If $\alpha=0$, then$|[x,y]|\leq b+1$ for any two vertices $x$ and $y$ of $\Gamma$ at distance 2.
Proof of Claim 1.
Let $M$ be the set of vertices consisting of neighbours of $x$ at distance at most 2 from $y$. Then
$|M|$ $=k-b_{2}=\beta r-(r-\frac{b^{2}-1}{b-1})(\beta-\alpha\frac{b^{2}-1}{b-1})$
$=(b+1)(\beta+r\alpha-(b+1)\alpha)=(b+1)(a_{1}+1-b\alpha).$
As every line has at least $a_{1}+2-(s-1)(c_{2}-1)$ vertices, every line of $[x,y]$ has at least $a_{1}+1-(s-1)(c_{2}-1)$ vertices in $M$, because $x\not\in M$.
Therefore
$|[x,y]|$ $\leq\frac{|M|}{a_{1}+1-(s-1)(c_{2}-1)}$ (8)
$=\frac{(b+1)(a_{1}+1-b\alpha)}{a_{1}+1-(s-1)(c_{2}-1)}=(b+1)(1+\frac{(s-1)(c_{2}-1)-b\alpha}{a_{1}+1-(s-1)(c_{2}-1)}).$ (9)
If $\alpha\geq 1$, by the SPLS(s) property we have $a_{1}+1\geq(2s-1)(c_{2}-1)+1=2(s-1)(c_{2}-1)+c_{2}$, and hence we have $|[x,y]|<2(b+1)$. If $\alpha=0$, then, by Condition (1) of the theorem, we have
$|[x,y]|\leq\frac{(b+1)\beta}{\beta-(s-1)b}<b+2.$
This shows the claim.
Claim 2
  1. (1)
    We have $|[x,y]|\leq b+1$ for any two vertices $x$ and $y$ of $\Gamma$ at distance 2;
  2. (2)
    If a vertex has distance 1 to a line, then it has at most $b+1$ neighbours on the line;
  3. (3)
    $\Gamma$ has the SPLS$(b+1,s)$ property.
Proof of Claim 2. By Claim 1, we may assume $\alpha\geq 1$. Let $q$ be the maximum integer such that there exist vertices $x,y$ at distance 2 with $|[x,y]|=q+1$. By Proposition 13 we have $q\geq b$.
We will first show that every line has at least $a_{1}+2-(s-1)q$ vertices. Let $\ell$ be a line and let $u\sim v$ be vertices on $\ell$. We’ll show that any line $\ell_{1}$ through $u$ different from $\ell$ contains at most $q$ neighbours of $v$ other than $u$. Let $w$ on $\ell_{1}$ at distance 2 from $v$. Each neighbour of $v$ other than $u$ on the line $\ell_{1}$ gives a unique line in $[v,w]$. This means that $\ell_{1}$ contains at most $q$ neighbours of $v$ besides $u$. As $u$ lies on at most $s$ lines, we obtain that $u$ and $v$ have at most $|\ell|-2+(s-1)q$ common neighbours. Hence $a_{1}\leq|\ell|-2+(s-1)q$, or equivalently, $|\ell|\geq a_{1}+2-(s-1)q$. Let $M$ be as in Claim 1. As every line has at least $a_{1}+2-(s-1)q$ vertices, every line of $[x,y]$ has at least $a_{1}+1-(s-1)q$ vertices in $M$. Therefore $(q+1)(a_{1}+1-(s-1)q)\leq|M|$. Let $f(x)=(x+1)(a_{1}+1-(s-1)x)$. Then $f(q)\leq|M|$. By Claim 1 we have $q\leq 2b$, and we have already seen $q\geq b$. We will show that $q=b$. Note that $|M|=(b+1)(a_{1}+1-b\alpha).$ The minimum of $f(x)$ for $b+1\leq x\leq 2b$ is the minimum of $f(b+1)$ and $f(2b)$.
First we show that $f(b+1)\leq f(2b)$ holds.
Now
$f(2b)-f(b+1)=(2b+1)(a_{1}+1-(s-1)2b)-(b+2)(a_{1}+1-(s-1)(b+1)).$
This implies that
$f(2b)-f(b+1)=(b-1)(a_{1}+1)-(s-1)(3b^{2}-b-2)=(b-1)(a_{1}+1-(s-1)(3b+2)).$
So $f(2b)\geq f(b+1)$ if $a_{1}+1\geq(s-1)(3b+2)$. We already found that
$a_{1}+1\geq 2(s-1)(c_{2}-1)+c_{2}=2(s-1)((b+1)(\alpha+1)-1)+(b+1)(\alpha+1).$
We see that $f(2b)\geq f(b+1)$ as $\alpha\geq 1$ and $b\geq 2$. Now we need to compare $f(b+1)$ with $|M|$. We find
$f(b+1)-|M|$ $=(b+2)(a_{1}+1-(s-1)(b+1))-(b+1)(a_{1}+1-b\alpha)$
$=a_{1}+1-(b+1)((b+2)(s-1)-b\alpha).$
As $a_{1}+1=\beta+\alpha(r-1)$ and $\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha$, we have
$f(b+1)>|M|.$
This means that $q\leq b$ and hence this shows Item (1) of this claim.
Let $x$ be a vertex and $\ell$ a line such that $d(x,\ell)=1$. Then $x$ has at most $c_{2}$ neighbours on $\ell$. This means that there exists a vertex $y$ on $\ell$ at distance 2 from $x$. There are at most $b+1$ lines in $[x,y]$ and each of them intersects in at most one vertex with $\ell$ and a line through $x$ that intersects with $\ell$ is in $[x,y]$. So $x$ has at most $b+1$ neighbours in $\ell$.
Now it follows that $\Gamma$ has the SPLS$(b+1,s)$ property, as through each vertex there at most $s$ lines. This shows the claim.
Now we will look at vertices at distance 3 in $\Gamma$.
Claim 3 Let $x,y$ be two vertices at distance 3 in $\Gamma$, and $\ell$ a line through $y$. Then the number of vertices $z$ at distance 2 from $x$ such that $z\in\ell$ is less than $(b+1)(b^{2}+b+1)$.
Proof of Claim 3. Let $U=\Gamma_{2}(x)\cap\ell$ and $W=\Gamma(x)\cap\Gamma_{2}(y)$. Then $|W|=c_{3}=(b^{2}+b+1)(1+\alpha(b+1)$. If $u\in U$, then $\Gamma(u)\cap\Gamma(x)\subseteq W$. Hence every vertex of $U$ has $c_{2}$ neighbours in $W$. Counting the number of edges between $U$ and $W$ we find $|U|c_{2}=\sum_{w\in W}|\Gamma(w)\cap\ell|.$ As a vertex outside $\ell$ has at most $b+1$ neighbours on $\ell$, we obtain $|U|c_{2}\leq|W|(b+1)=(b+1)(b^{2}+b+1)(1+\alpha(b+1))$. Now as $c_{2}=(b+1)(\alpha+1)$, we see that $|U|\leq\frac{(b+1)(b^{2}+b+1)(1+\alpha(b+1))}{(b+1)(\alpha+1)}<(b^{2}+b+1)(b+1)$, and this shows the claim.
Claim 4 If $\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$, then for any pair of vertices $x,y$ at distance 2, the number $|[x,y]|$ satisfies $|[x,y]|=b+1$.
Proof of Claim 4. As above let $M$ be the set of neighbours of $x$ that have distance 1 or 2 from $y$. Then $|M|=k-b_{2}=(b+1)\beta+\alpha(b+1)r-\alpha(b+1)^{2}$. Let $c:=(b+1)(b^{2}+b+1).$ As $\Gamma$ has the SPLS$(b+1,s)$ propert, by Claim 1, we only need to show that $|[x,y]|\geq b+1$. Assume that $|[x,y]|\leq b$. Any line through $x$ contains at most $\beta+1$ vertices, and therefore, there are at most $\beta$ vertices in $M$.
A line through $x$ that is not in $[x,y]$ contains a vertex at distance 3 from $y$, and hence by Claim 3 contains at most $c-2$ vertices in $M$. As $x$ is on at most $s$ lines, it follows that $|M|\leq|[x,y]|\beta+(s-|[x,y]|)(c-2)\leq b\beta+(s-b)(c-2)$, as $c-2<\beta$. As $|M|=k-b_{2}=(b+1)\beta+\alpha(b+1)r-\alpha(b+1)^{2}$, we obtain $\beta\leq(s-b)(c-2)-\alpha(b+1)r+\alpha(b+1)^{2}$. So, if $\beta>(s-b)(c-2)-\alpha(b+1)r+\alpha(b+1)^{2}=((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$, then $|[x,y]|=b+1$, as we have $s\geq r\geq b^{2}+b+1$ and $\alpha\leq b$. This shows Claim 4.
Claim 5 If $\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b$, then for each pair of a vertex $x$ and a line $\ell$, with $d(x,\ell)=1$ and vertices $y_{1},y_{2}$ on $\ell$ with $d(x,y_{1})=d(x,y_{2})=2$, we have that the sets of lines $[x,y_{1}]$ and $[x,y_{2}]$ are equal.
Proof of Claim 5. Assume that there are vertices $y_{1},y_{2}\in\Gamma_{2}(x)$ that are both on $\ell$ such that $[x,y_{1}]\neq[x,y_{2}]$. Put $S_{i}:=\Gamma_{i}(x)\cap\ell$ and $n_{i}:=|S_{i}|$ for $i=1,2.$ Further, put $T:=[x,y_{1}]\cap[x,y_{2}]$ and $t:=|T|$. Then $1\leq n_{1}\leq t\leq b$, as two distinct lines intersect in at most one vertex, $[x,y_{1}]\neq[x,y_{2}]$ and $|[x,y_{1}]|=|[x,y_{2}]|=b+1$. Let $T_{2}$ be the set of lines through $x$ that are not in $T$. Then $t_{2}:=|T_{2}|\leq s-t$, since $x$ lies on at most $s$ lines. Let $\ell_{2}\in T_{2}$ be a line through $x$ that is not in $[x,y_{1}]$. Then there exists a vertex $u$ on $\ell_{2}$ at distance 3 from $y_{1}$. Now we obtain $|\Gamma_{2}(u)\cap\ell|\leq c-1$ where $c=(b+1)(b^{2}+b+1)$. As $u\sim x$ and $d(u,y_{1})=3$, we have $S_{1}\subseteq\Gamma_{2}(u)\cap\ell$. As a consequence we find $|\Gamma_{2}(u)\cap S_{2}|\leq|\Gamma_{2}(u)\cap\ell|-|S_{1}|\leq c-1-n_{1}.$ It follows that there are at most $c-1-n_{1}$ vertices $v\in S_{2}$ with $\ell_{2}\in[x,v]$, as $d(u,v)\leq 2$ for such a vertex $v$.
On the other hand, if $v\in S_{2}$, then $[x,v]$ contains at least $b+1-t$ lines of $T_{2}$. By counting pairs $(v,\ell_{2})$ with $v\in S_{2}$ and $\ell_{2}\in T_{2}\cap[x,v]$ we obtain
$n_{2}(b+1-t)\leq|T_{2}|(c-1-n_{1})\leq(s-t)(c-1-n_{1}).$
As $t\leq b$ and $s\geq r>b+1$, we obtain $(b-t)(s-b-1)\geq 0$ which is equivalent with $s-t\leq(b+1-t)(s-b)$. As a consequence we have $n_{2}\leq(s-b)(c-1-n_{1})$. This means that $|\ell|=n_{1}+n_{2}\leq n_{1}+(s-b)(c-1-n_{1})\leq 1+(s-b)(c-2)$, as $n_{1}\geq 1$ and $s\geq b+2$. By Claim 2, the graph $\Gamma$ has the SPLS$(b+1,s)$ property, and hence we have $|\ell|\geq a_{1}+2-(s-1)b$. Hence $\beta+\alpha(r-1)-(s-1)b+1=a_{1}+2-(s-1)b\leq 1+(s-b)(c-2)$ and therefore $\beta\leq(s-b)(c-2+b)-\alpha(r-1)+b^{2}-b$. So, if $\beta>(s-b)(c-2+b)-\alpha(r-1)+b^{2}-b$, the sets $[x,y_{1}]$ and $[x,y_{2}]$ are equal and this shows Claim 5.
Claim 6 Assume that $x$ is a Delsarte vertex and that $y$ is a neighbour of $x$. Then $y$ is also a Delsarte vertex.
Proof of Claim 6. Let $C_{1},C_{2},\ldots,C_{r}$ be the lines through $x$ and without loss of generality, assume $y$ in $C_{1}$. The $C_{i}$’s are all Delsarte cliques by the fact that $x$ is a Delsarte vertex. Let $C_{1},D_{2},\ldots,D_{\sigma}$ be the lines through $y$. Then $\sigma\geq r$, and if equality holds, then $y$ is a Delsarte vertex. Let $\tilde{C}_{i}$ be the set of vertices in $C_{i}$ that are not neighbours of $y$, and $\tilde{D}_{i}$ be the set of vertices in $D_{i}$ that are not neighbours of $x$ for $i\geq 2$. Let $\Delta$ be the bipartite graph with vertex set $\{\tilde{C}_{2},\tilde{C}_{3},\ldots,\tilde{C}_{r},\tilde{D}_{2},\tilde{D}_{3},\ldots,\tilde{D}_{\sigma}\}$, and $\tilde{C}_{i}$ is adjacent to $\tilde{D}_{j}$ if and only if there are vertices $u\in\tilde{C}_{i}$ and $v\in\tilde{D}_{j}$ that are adjacent, for $i,j\geq 2$. We will obtain two bounds on the number of edges of $\Delta$ by double-counting. Let $v_{1},v_{2}$ be two distinct vertices of $\tilde{D}_{j}$ for some $j\geq 2$. By Claims 4 and 5, we have $[x,v_{1}]=[x,v_{2}]$ both with cardinality $b+1$. As ${C}_{1}\in[x,v_{1}]$, we have that the valency $k_{j}$ of $\tilde{D}_{j}$ is at most $b$. Let $C_{i}\in[x,v_{1}]$ for some $i\geq 2$. Then, as $C_{i}$ is a Delsarte clique, there are $\alpha+1$ neighbours of $v_{1}$ inside $C_{i},$ by Lemma 7. Note that $|\Gamma(y)\cap C_{i}-\{x\}|=\alpha$, therefore $\Gamma(v_{1})\cap\tilde{C_{i}}\neq\emptyset$. This means that $k_{j}$ is equal to $b$ and the number $\varepsilon$ of edges of $\Delta$ satisfies $\varepsilon=b(\sigma-1).$ Let $u_{1},u_{2}$ be two distinct vertices of $\tilde{C}_{i}$ for some $i\geq 2$. Again by Claims 4 and 5, we have $[y,u_{1}]=[y,u_{2}]$ both with cardinality $b+1$. As ${C}_{1}\in[y,u_{1}]$ we have that the valency $\kappa_{i}$ of $\tilde{C}_{i}$ is at most $b$. Therefore $\varepsilon\leq(r-1)b$ and hence $\sigma\leq r$. This means that $y$ is also a Delsarte clique.
By Proposition 13 we are done as we have at least one Delsarte vertex and hence all the vertices must be Delsarte vertices. This shows that $\Gamma$ is geometric.
Now we can calculate the $\phi_{j}$’s and $\tau_{j}$’s of Equation (3) for a geometric distance-regular graph with classical parameters.
Lemma 16
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$, and assume that $\Gamma$ is geometric. Then the numbers $\phi_{j}$’s and $\tau_{j}$’s of Equation (3) satisfy
$\phi_{j}=1+\alpha[j],\ \ (j=0,1,\ldots,D-1),\ \ \tau_{j}=[j]\ \ (j=1,2,\ldots,D),$
where $[j]=\frac{b^{j}-1}{b-1}$.
Proof. The expressions for the $\phi_{j}$’s follow from Lemma 7 and for the $\tau_{j}$’s follow now from Equation (3).
Now we first consider $\alpha=0$.
Proposition 17
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 1$, $\alpha=0$, $\beta>1$ and $D\geq 3$. If $\Gamma$ is geometric, then $\Gamma$ is a dual polar graph or a Hamming graph.
Proof. As $\alpha=0$ we have $a_{1}=\beta-1>0$ and $\Gamma$ is locally the disjoint union of cliques of order $\beta$. By [1, Theorem 9.4.4], we are now done.
For geometric distance-regular graphs with classical parameters, we can provide a better bound for $\beta$ than the one in Claim 5 of Theorem 15, using the same discussion as before.
Theorem 18
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $r=[D]$. If $\beta>(r-\alpha-1)\alpha b+\alpha$, then for each pair of a vertex $x$ and a line $\ell$, with $d(x,\ell)=1$ and vertices $y_{1},y_{2}$ on $\ell$ with $d(x,y_{1})=d(x,y_{2})=2$, we have that the sets of lines $[x,y_{1}]$ and $[x,y_{2}]$ are equal.
Proof. Let $\Gamma$ be geometric with respect to $\mathcal{C}$. Let $x,y$ be two vertices of $\Gamma$ at distance 2. We will first show $|[x,y]|=b+1$. If $\ell$ is a line in $[x,y]$, then $d(y,\ell)=1$ as $\ell$ is a Delsarte clique, and hence $|[x,y]|=\tau_{2}=b+1.$ Let $u,v$ be two vertices at distance 3 in $\Gamma$, and $m$ a line through $u$ and at distance 2 from $v$. Then the number of vertices $z$ at distance 2 from $v$ such that $z\in m$ is equal to $\phi_{2}=1+\alpha(b+1)$ and hence at most $b^{2}+b+1$.
Let $\ell$ be a line at distance 1 from $x$. Assume that there are vertices $y_{1},y_{2}$ at distance 2 from $x$ that are both on $\ell$ such that $[x,y_{1}]\neq[x,y_{2}]$. Put, for $i=1,2$, $S_{i}:=\Gamma_{i}(x)\cap\ell$ and $n_{i}:=|S_{i}|$. By Lemma 7 we have $|\ell|=\beta+1$, $n_{1}=\phi_{1}=\alpha+1$ and $n_{2}=|\ell|-n_{1}=\beta-\alpha$, as $\Gamma$ is geometric. This means that there are $\alpha+1$ lines through $x$ that intersect $\ell$, as two distinct lines intersect in at most one point and $n_{1}=\alpha+1$. Those lines are in $T:=[x,y_{1}]\cap[x,y_{2}]$. Let $t:=|T|$. As $|[x,y_{1}]|=|[x,y_{2}]|=b+1$, and $[x,y_{1}]\neq[x,y_{2}]$, we have $\alpha+1=n_{1}\leq t\leq b$. As there are exactly $r$ lines through $x$, there are exactly $r-t$ lines through $x$ that are not in $T$. Let $u\sim x$ such that $d(y_{1},u)=3$. Then the line $\ell_{1}$ through both $u$ and $x$ does not lie in $[x,y_{1}]$. Then $d(u,\ell)=2$ and hence $|\Gamma_{2}(u)\cap\ell|=\phi_{2}=1+\alpha(b+1)$, by Lemma 7. As $u\sim x$ and $d(u,y_{1})=3$, we find that every vertex in $S_{1}$ is at distance 2 from $u$. As a consequence $|\Gamma_{2}(u)\cap S_{2}|\leq|\Gamma_{2}(u)\cap\ell|-|S_{1}|=1+\alpha(b+1)-(1+\alpha)=\alpha b$. It follows that there are at most $\alpha b$ vertices $v$ in $S_{2}$ such that $\ell_{1}\in[x,v]$, as this implies $d(u,v)\leq 2$.
On the other hand, for $v\in S_{2}$, the set $[x,v]$ contains at least $b+1-t$ lines through $x$ that are not in $T$. By counting pairs $(v,\ell_{2})$ with $v\in S_{2}$ and lines $\ell_{2}$ in $[x,v]$ that are not in $T$ we have $\beta-\alpha\leq(\beta-\alpha)(b+1-t)=n_{2}(b+1-t)\leq(r-t)\alpha b$, as $\ell_{2}$ contains a vertex $w$ that is in $\Gamma_{3}(y_{1})\cup\Gamma_{3}(y_{2})$ and $t\leq b$.
Therefore $\beta-\alpha\leq(r-t)\alpha b$. So, if $\beta>(r-\alpha-1)\alpha b+\alpha$, the sets $[x,y_{1}]$ and $[x,y_{2}]$ are equal. This shows the theorem.
In the next theorem we give a bound on $\beta$ to insure that a distance-regular graph with classical parameters is geometric.
Theorem 19
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $r=[D]$. If
$\beta\geq\max\{\frac{2b+4}{2b+3}r(b+2)(\alpha b+b+\alpha),\frac{2b+4}{2b+3}r((b+1)(b^{2}+b+2)-3)\},$
then $\Gamma$ is geometric.
Proof. We only need to show the existence a positive integer $s$ such that the conditions in both Theorem 9 and Theorem 15 are satisfied. As we already have seen in the proof of Theorem 14, Equations (4) and (5), the conditions in Theorem 9 can be written as
$\beta>\frac{s+1}{2(s+1-r)}\left((\alpha b+\alpha+b)s-2\alpha(r-1)\right),$ (10)
and
$\beta>(\alpha b+b+\alpha)2s-\alpha r-(\alpha+1)b.$ (11)
Let $s_{0}:=\frac{2b+4}{2b+3}r$ be a rational number and $s:=\lfloor s_{0}\rfloor$. Then $\frac{s+1}{s+1-r}<\frac{s_{0}}{s_{0}-r}=2(b+2).$ So to show that Conditions (10) and (11) hold, it suffices to show that the following two conditions hold:
$\beta>(b+2)\left((\alpha b+b+\alpha)s_{0}-2\alpha(r-1)\right),$ (12)
and
$\beta$ $>2(\alpha b+b+\alpha)s_{0}-\alpha r-(\alpha+1)b.$ (13)
But these two conditions are satisfied as
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
This shows that $\Gamma$ satisfies the SPLS$(s)$ property. Now we need to consider the conditions in Theorem 15. These are
$\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha,$ (14)
$\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2},$ (15)
and
$\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b.$ (16)
As
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}$
and $s_{0}=\frac{2b+4}{2b+3}r$, we find that
$\beta>(b+2)(b+1)(s_{0}-1)-(b(b+1)+r-1)\alpha,$ (17)
$\beta>((b+1)(b^{2}+b+1)-2)(s_{0}-b)-\alpha(b+1)r+\alpha(b+1)^{2},$ (18)
and
$\beta>((b+1)(b^{2}+b+1)-2+b)(s_{0}-b)-\alpha(r-1)+b^{2}-b$ (19)
are satisfied and hence the conditions of Theorem 15 are satisfied. This shows that $\Gamma$ is geometric and this finishes the proof.
As a corollary we have the following consequence for distance-regular graph with classical parameters $(D,b,0,\beta)$.
Corollary 20
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$, such that $b\geq 2$, $\alpha=0$ and $D\geq 3$. Let $r=[D]$. Then $\beta<2\frac{b+2}{2b+3}(b^{3}+2b^{2}+3b-1)r$ holds.
Proof. For $\beta=1$ the corollary is clearly true, so we may assume $\beta>1$. If $\beta\geq 2\frac{b+2}{2b+3}(b^{3}+2b^{2}+3b-1)r$, then $\Gamma$ is geometric by Theorem 19. By Proposition 17, we find that $\Gamma$ is a dual polar graph or a Hamming graph. In the first case we have $\beta\leq b^{2}$, by [1, Table 6.1]. In the last case we have $b=1$. So the theorem follows.
5 Dual Pasch Axiom
In [1, Theorem 9.3.9], they give sufficient conditions to assure that the point graph of a partial linear space is a Grassmann graph. This covers the case $\alpha=b$ for distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ when $b=\alpha\geq 2$, $D\geq 3$ and $\beta$ large enough. Therefore we will assume $11\leq\alpha\leq b-$ 1 from now on.
Definition 21
Assume that $\Gamma$ is a distance-regular graph such that it is the point graph of a partial linear space $X=(V(\Gamma),\mathcal{L},\in)$. Then $\Gamma$ satisfies the dual Pasch axiom if for each pair of two adjacent vertices $x,y$ of $\Gamma$, the set of common neighbours of $x$ and $y$, outside the line through $x$ and $y$, forms a clique.
We will show that a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$ satisfies the dual Pasch axiom if $\beta$ is large enough.
Lemma 22
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Let $\mathcal{C}$ be a set of Delsarte cliques of $\Gamma$ such that $\Gamma$ is the point graph of the partial linear space $X=(V(\Gamma),{\mathcal{C}},\in)$. Let $r=[D]$. If $\beta>(r-\alpha-1)\alpha b+\alpha$, then $\Gamma$ satisfies the dual Pasch axiom.
Proof. Note that $\alpha$ is an integer and $r\geq b^{2}+b+1$, as $D\geq 3$. For each pair of a vertex $x$ and a line $\ell$ with $d(x,\ell)=1$, and vertices $y_{1},y_{2}$ on $\ell$ with $d(x,y_{1})=d(x,y_{2})=2$, we have that the sets of lines $[x,y_{1}]$ and $[x,y_{2}]$ are equal, by Theorem 18. Let $x$ and $y$ be two adjacent vertices and let $\ell$ be the line through $x$ and $y$. Let $U$ be the set of common neighbours of $x$ and $y$ that are not on $\ell$. We will show that two distinct vertices in $U$ are adjacent. As $\ell$ contains exactly $\beta+1$ vertices and $a_{1}=\beta-1+(r-1)\alpha$, we find that $|U|=a_{1}-\beta+1=(r-1)\alpha$. Let $u\in U$. We will first show that $u$ has at least $|U|-1-\alpha(b-1)=(r-b)\alpha-1$ neighbours in $U$. Let $m$ be the (unique) line through $x$ and $u$. Let $W\subseteq U$ be the set of vertices in $U$, distinct from $u$, that are not adjacent to $u$. Then $|W|\leq|U|-\alpha=(r-2)\alpha$, as there are exactly $\alpha$ neighbours of $y$ besides $x$ on $m$. As any vertex $w$ of $W$ has exactly $\alpha$ neighbours on $m$ besides $x$, the vertices in $W$ have in total at most $(r-2)\alpha^{2}$ neighbours on $m$ besides $x$. Note that $y$ has exactly $\alpha$ neighbours on $m$ besides $x$. As $b-1\geq\alpha\geq 1$ and $b\geq 2$, we have $\beta>(r-\alpha-1)\alpha b+\alpha>(r-2)\alpha^{2}+\alpha$. As $m$ has exactly $\beta+1$ vertices and $\beta>(r-2)\alpha^{2}+\alpha$, there exists a vertex $p$ on $m$ that has distance 2 to all vertices in $W$ and is non-adjacent to $y$.
Let $w\in W$. Consider the (unique) line $h$ through $y$ and $w$. We will show that $h\in[y,p]$. Note that we have $h\in[w,u]$. As $u$ and $p$ are both on $m$ and $d(u,w)=d(p,w)=2$, we have that the sets $[w,p]$ and $[w,u]$ are equal, by Theorem 18. Therefore we have $h\in[w,p]$. This means that there exists a vertex $q$ on $h$ at distance 1 from $p$. This shows that $h\in[y,p]$.
Now we can give a better upper bound for $|W|$. There are exactly $b-1$ lines in $[y,p]$ besides $\ell$ and the unique line through $u$ and $y$, and each of those lines contains exactly $\alpha$ neighbours of $x$ besides $y$, we find $|W|\leq(b-1)\alpha$. This means that $u$ has at least $|U|-1-\alpha(b-1)=(r-b)\alpha-1$ neighbours in $U$. But this implies that two non-adjacent distinct vertices $u$ and $w$ in $U$ have at least $(r-2b+1)\alpha$ common neighbours in $U$. On the other hand they have exactly $c_{2}=(\alpha+1)(b+1)$ common neighbours.
The vertex $u$ has at least one neighbour on $h$ that is not adjacent to $x$ and similar $w$ has at least one neighbour on $m$ that is not adjacent to $y$, and $x,y$ are common neighbours of $u$ and $w$. So we obtain $4+(b^{2}-b+2)\alpha\leq 2+2+(r-2b+1)\alpha\leq(\alpha+1)(b+1)$ as $r\geq b^{2}+b+1$. But this is impossible. This means that all pairs of distinct vertices in $U$ are adjacent and this shows the lemma.
Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$ and with classical parameters $(D,b,\alpha,\beta)$ with $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$ satisfying the dual Pasch axiom. A maximal clique that is not in $\mathcal{C}$ is called an assembly.
Proposition 23
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Let $\mathcal{C}$ be a set of Delsarte cliques of $\Gamma$ such that $\Gamma$ is the point graph of the partial linear space $X=(V(\Gamma),{\mathcal{C}},\in)$. Let $r=[D]$. Assume that $\Gamma$ satisfies the dual Pasch axiom. Then each pair of two adjacent vertices $x$ and $y$ lie in a unique assembly, say $M_{x,y}$, and $M_{x,y}$ has order $\alpha r+1$ and therefore $\beta\geq\alpha r$. As a consequence, we have that each vertex lies in exactly $\frac{\beta}{\alpha}$ assemblies.
Proof. Let $U$ be the set of common neighbours of $x$ and $y$ that are not on the unique line $\ell$ through $x$ and $y$. As $a_{1}=\beta-1+\alpha(r-1)$ and $\ell$ contains exactly $\beta+1$ vertices, it follows that $|U|=\alpha(r-1)$. By the dual Pasch axiom, every two distinct vertices in $U$ are adjacent. Let $u\in U$. Then $u$ has exactly $1+\alpha$ neighours in $\ell$, as $u$ is not on $\ell$. If $\alpha=1$, then the induced subgraph $C$ on $U\cup\{x,y\}$ is a maximal clique of $\Gamma$ as $C$ intersects $\ell$ in exactly $\alpha+1=2$ vertices. This shows that $x,y$ lie in a unique assembly $M_{x,y}$ and that $M_{x,y}$ has order $\alpha(r-1)+\alpha+1=\alpha r+1$ when $\alpha=1$.
From now we assume $\alpha\geq 2$. Let $z$ be a neighbour of $u$ in $l$ distinct from $x$ and $y$, and let $v\in U$ be distinct from $u$. Consider the line $\ell_{1}$ through $x$ and $u$, and the line $\ell_{2}$ through $y$ and $u$. Note that $z\notin\ell_{1}\cup\ell_{2}$, as two lines do not intersect in more than one vertex. The vertex $v$ cannot lie on both lines of $\ell_{1}$ and $\ell_{2}$, so we may assume without loss of generality that $v\notin\ell_{1}$. The vertices $z$ and $v$ are common neighbours of $x$ and $u$, both not lying on the line $\ell_{1}$, and they are distinct since $v$ is not on $\ell$. By the dual Pasch axiom, the vertices $v$ and $z$ are adjacent. This shows that the subgraph $C$ induced on the set $U\cup\{z\mbox{ on }\ell\mid z\sim u\}$ is a clique and $C$ is maximal, as $C$ intersects $\ell$ in exactly $\alpha+1$ vertices. Hence $C$ is the maximal clique containing $\{x,y\}\cup U$. This shows that $x,y$ lie in a unique assembly $M_{x,y}$ and that $M_{x,y}$ has order $\alpha(r-1)+\alpha+1=\alpha r+1$, when $\alpha\geq 2$. So we have shown that each assembly has exactly $\alpha r+1$ vertices. This implies that $\beta\geq\alpha r$ as a clique has at most $\beta+1$ vertices. Therefore each vertex $x$ lies in exactly $\frac{k}{\alpha r}=\frac{\beta}{\alpha}$ assemblies, as $k=r\beta$. This shows the proposition.
Remark 3
  1. (1)
    Note that the Grassmann graph $J_{q}(2D,D)$ where $q$ is a prime power and $D\geq 2$ has $\alpha r=\beta$.
  2. (2)
    Jurišić and Vidali [8] showed that for a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $b\geq 2$ and $D\geq 3$, $\beta\geq 1+\alpha[D-1]$ holds, and the equality case is not possible for $D\geq 10$ by Koolen et al., [9, Corollary 7.3] and [8, Proposition 2].
The above proposition shows that through any edge, there is a unique assembly, and that all the assemblies have order $\alpha r+1$. Now we show that any local graph is an $\alpha$-clique extension of a grid.
Theorem 24
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Let $\mathcal{C}$ be a set of Delsarte cliques of $\Gamma$ such that $\Gamma$ is the point graph of the partial linear space $X=(V(\Gamma),{\mathcal{C}},\in)$. Let $r=[D]$. For a vertex $x$, let $\Delta_{x}$ be the local graph at vertex $x$. Assume that $\Gamma$ satisfies the dual Pasch axiom. Let $\ell$ be a line, $M$ an assembly and $u$ be a vertex of $\Gamma$ not in $M$. Then the following conditions hold:
  1. (1)
    If $\ell$ intersects with $M$, then they intersect in exactly $\alpha+1$ vertices;
  2. (2)
    If $u$ has a neighbour in $M$, then $u$ has exactly $\alpha+1$ neighbours in $M$ and there is exactly one line through $u$ that intersects $M$.
As a consequence we have that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid.
Proof. Let $M$ be an assembly and $\ell$ be a line in $\mathcal{C}$ and assume that $M$ and $\ell$ intersect in a vertex $w$. Let $\ell_{1}:=\ell,\ell_{2},\ldots,\ell_{r}$ be the lines through $w$. As $M-\ell_{i}\neq\emptyset$ for $i=1,2,\ldots,r$ and each vertex outside $\ell_{i}$ has at most $\alpha+1$ neighbours in $\ell_{i}$, we have $|M\cap\ell_{i}|\leq\alpha+1$ for $i=1,2,\ldots,r$. Note that $M-\{w\}=\bigcup\limits_{i=1}^{r}(M\cap\ell_{i}-\{w\})$. Therefore $\alpha r=|M-\{w\}|\leq\sum\limits_{i=1}^{r}(|M\cap\ell_{i}|-1)\leq\sum\limits_{i=1}^{r}\alpha=\alpha r$, and hence $|M\cap\ell_{i}|=\alpha+1$ for $1\leq i\leq r$. In particular, $|M\cap\ell|=\alpha+1$. This shows Item (1) of the theorem.
Now let $u$ have a neighbour $v$ in $M$. Let $m$ be the line through $u$ and $v$. Then $|m\cap M|=\alpha+1$. Let $w\in M$ such that $w$ is not on $m$. Then $w$ has exactly $\alpha+1$ neighbours on $m$, and those must be in $M$, since $|m\cap M|=\alpha+1$. This shows that $u$ has exactly $\alpha+1$ neighbours in $M$ and that there is a unique line through $u$ that intersects $M$. This shows Item (2) of the theorem.
Let $x$ be a vertex of $\Gamma$. Let $\ell_{1},\ell_{2},\ldots,\ell_{r}$ be the lines through $x$ and $M_{1},M_{2},\ldots,M_{t}$, where $t=\frac{\beta}{\alpha}$, be the assemblies through $x$. Let $1\leq i,a,\leq r$ and $1\leq j,b\leq t$. Then $|\ell_{i}\cap M_{j}|=\alpha+1$ and $\ell_{i}\cap M_{j}\cap\ell_{a}\cap M_{b}=\emptyset$ unless $(i,j)=(a,b)$. This shows that $\Delta_{x}$ is the $\alpha$-clique extension of the $\frac{\beta}{\alpha}\times r$-grid.
Definition 25
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Assume that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. Assume $\beta>\alpha r$. Let $M$ be an assembly. For a vertex $x$ at distance $i$ from $M$, define $\phi^{*}_{i}(M,x)$ as the number of vertices in $M$ at distance $i$ from $x$ for $i=0,1,\ldots,D$. For two vertices $x,y$ at distance $j$ $(j=1,2,\ldots,D)$, let $\tau^{*}_{j}(x,y)$ be the number of assemblies that contain $y$ and are at distance $j-1$ from $x$.
Now we will show that $\phi^{*}_{i}(M,x)$ does not depend on either $M$ or $x$, but only on $i$. Also we will show that $\tau_{j}^{*}(x,y)$ does not depend on the pair $(x,y)$, only on $j$.
Theorem 26
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Assume that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. Assume $\beta>\alpha r$. Let $M$ be an assembly and $x$ be a vertex at distance $i$ from $M$ for $i=0,1,\ldots,D$. Then $\phi_{i}^{*}(M,x)=\phi^{*}_{i}=1+\alpha[i]$ holds. Let $x$ and $y$ be two vertices at distance $j$ where $j=1,2,\ldots,D$. Then $\tau^{*}_{j}(x,y)=\tau^{*}_{j}=[j]$ holds. Let $x$ and $y$ be two vertices at distance $h$ where $h=0,1,\cdots,D-1$. Then the induced subgraph on the set $B_{h}(x,y)$ is the $\alpha$-clique extension of a $(\frac{\beta}{\alpha}-[h])\times(r-[h])$-grid.
Proof. We will show the theorem by induction on $i$. The assumption that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid shows the theorem for $i=0$.
Let $i\geq 1$ and assume that the theorem holds for $10\leq t\leq i-$ 1. Let $x$ and $y$ be two vertices at distance $i$. Let $M$ be an assembly through $y$ at distance $i-1$ from $x$. As $\phi^{*}_{i-1}(M,x)=1+\alpha[i-1]$ and $c_{i}=[i](1+\alpha[i-1])$, there are exactly $[i]$ assemblies through $y$ at distance $i-1$ from $x$. This implies that $\tau_{i}^{*}(x,y)=[i]$. As this holds for any pair of vertices $x,y$ at distance $i$, we find $\tau^{*}_{i}=[i]$.
Now we will show that the induced subgraph on $B_{i}(x,y)$ is the $\alpha$-clique extension of the $(\frac{\beta}{\alpha}-[i])\times(r-[i])$-grid. Note that for all $u\in C_{i}(x,y)$ and $v\in B_{i}(x,y)$ the distance between $u$ and $v$ is 2. We also have $b_{i}=(\beta-\alpha[i])(r-[i])$.
By Lemma 16, there are exactly $[i]$ lines through $y$ at distance $i-1$ from $x$. So there are exactly $r-[i]$ lines through $y$ at distance $i$ from $x$. As $\tau^{*}_{i}=[i]$, there are exactly $[i]$ assemblies through $y$ at distance $i-1$ from $x$. Again there are exactly $\frac{\beta}{\alpha}-[i]$ assemblies through $y$ at distance $i$ from $x$. As every vertex in $B_{i}(x,y)$ is at distance 2 from $C_{i}(x,y)$, we see that the induced subgraph on $B_{i}(x,y)$ is an induced subgraph of the $\alpha$-clique extension of a $(r-[i])\times(\frac{\beta}{\alpha}-[i])$-grid. As $b_{i}=(r-[i])(\beta-\alpha[i])$, we find that the induced subgraph on $B_{i}(x,y)$ is the $\alpha$-clique extension of a $(r-[i])\times(\frac{\beta}{\alpha}-[i])$-grid.
Let $M$ be an assembly through $y$ at distance $i$ from $x$. Then $M$ has exactly $(r-[i])\alpha$ vertices at distance $i+1$ from $x$ and hence exactly $1+\alpha[i]$ vertices at distance $i$ from $x$. This shows that $\phi^{*}_{i}(M,x)=1+\alpha[i]=\phi^{*}_{i}$.
This completes the proof.
Theorem 27
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Assume that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. Assume $\beta>\alpha r$. For two vertices $x,y$ at distance $i$, $i=2,3,\ldots,D$, the induced subgraph on the set $C_{i}(x,y)$ is an induced subgraph of the $[i]\times[i]$-grid.
Proof. We first show the following claim.
Claim 1 Let $\ell$ be a line and $x$ a vertex at distance $i$ from $\ell$, for $i=1,2,\ldots,D-1$. Then for a pair of distinct vertices $y_{1},y_{2}\in\Gamma_{i}(x)\cap\ell$, we have $|\Gamma(y_{1})\cap\Gamma(y_{2})\cap\Gamma_{i-1}(x)|=1+\alpha[i-1]$ holds.
Proof of Claim 1. Let $T:=\Gamma_{i}(x)\cap\ell$ and $S:=\{u\mid u\in\Gamma_{i-1}(x)\mbox{ and }d(u,\ell)=1\}$. By double -counting the pairs $(u,v)$ of adjacent vertices where $u\in S$ and $v\in T$, we obtain $|T|c_{i}=|S|(\alpha+1)$. By Lemma 16, we have $|T|=\phi_{i}=1+\alpha[i]$, and as a consequence
$|S|=\frac{1+\alpha[i]}{\alpha+1}c_{i}=\frac{1+\alpha[i]}{\alpha+1}(1+\alpha[i-1])[i].$
For each pair of distinct vertices $v_{1},v_{2}\in T$, the set $\Gamma(v_{1})\cap\Gamma(v_{2})\cap\Gamma_{i-1}(x)\subseteq M_{v_{1},v_{2}}\cap\Gamma_{i-1}(x)$, where $M_{v_{1},v_{2}}$ is the unique assembly through $v_{1}$ and $v_{2}$, has at most $\phi^{*}_{i-1}=1+\alpha[i-1]$, by Theorem 26. Let $R:=\{(u,\{v_{1},v_{2}\})\mid u\in S\mbox{ and }v_{1},v_{2}\in T\cap\Gamma(u),v_{1}\neq v_{2}\}$. By double-counting we obtain $|S|{\alpha+1\choose 2}=|R|\leq{|T|\choose 2}(1+\alpha[i-1]).$ Using $|T|=1+\alpha[i]$ and $|S|=\frac{1+\alpha[i]}{\alpha+1}c_{i}=\frac{1+\alpha[i]}{\alpha+1}(1+\alpha[i-1])[i]$ we see that $|R|={|T|\choose 2}(1+\alpha[i-1]).$ This means that $|\Gamma(v_{1})\cap\Gamma(v_{2})\cap\Gamma_{i-1}(x)|=|M_{v_{1},v_{2}}\cap\Gamma_{i-1}(x)|=1+\alpha[i-1]$. This shows the claim.
By Lemma 16 and Theorem 26, we have $\tau_{i}^{*}=\tau_{i}=[i]$ for $i=1,2,\ldots,D$. To show Theorem 27, we just need to show that for any pair of vertices $x$ and $y$ at distance $i$ $(i=2,3,\ldots,D)$ and any assembly $M$ and line $\ell$ through $y$, $|\ell\cap M\cap\Gamma_{i-1}(x)|\leq 1$ holds. We will show this by contradiction. So, assume there are two distinct vertices $z_{1},z_{2}\in\ell\cap M\cap\Gamma_{i-1}$. Then, by Claim 1, we have $|\Gamma(z_{1})\cap\Gamma(z_{2})\cap\Gamma_{i-2}(x)|=1+\alpha[i-2]\geq 1$. Let $u\in\Gamma(z_{1})\cap\Gamma(z_{2})\cap\Gamma_{i-2}(x)$. Then the vertex $u$ should lie on the unique assembly containing $z_{1}$ and $z_{2}$, or the unique line containing $z_{1}$ and $z_{2}$, and both of them contain $y$. This is impossible as $d(u,y)=2$. This is a contradiction, which proves the theorem.
Now we show the existence of a 2-$(\alpha(b+1)+1,\alpha+1,1)$-design.
Theorem 28
Let $\Gamma$ be a geometric distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$, $b-1\geq\alpha\geq 1$ and $D\geq 3$. Assume that for any vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. Assume $\beta>\alpha r$. Let $M$ be an assembly and $x$ be a vertex at distance 2 from $M$ Let $B=\Gamma_{2}(x)\cap M$. Let $L(B)=\{\ell\cap B\mid\ell\mbox{ is a line such that }|\ell\cap B|\geq 2\}$. Then $(B,L(B))$ is a 2-$(\alpha(b+1)+1,\alpha+1,1)$-design. In particular $\alpha+1$ divides $b(b+1)$.
Proof. Let $M$ be an assembly and $x$ a vertex at distance 2 from $M$. Let $B=\Gamma_{2}(x)\cap M$. We first show the following claim.
Claim 1 For each line $\ell$ with $\ell\cap M\neq\emptyset,$ $|\ell\cap B|\geq 2$ holds if and only if $\ell\cap\Gamma(x)\neq\emptyset$. In particular, if $|\ell\cap B|\geq 2$, then $\ell\cap M=\ell\cap B$.
Proof of Claim 1. Let $\ell$ be a line and $y,z\in\ell\cap B$ be two distinct vertices. We will consider the local graph $\Delta_{y}$ at vertex $y$. By assumption, the graph $\Delta_{y}$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. As $d(x,M)=2$, there exists a vertex $u\in M$ at distance 3 from $x$. By Theorem 26, the induced subgraph on the set $B_{2}(x,y)$ is the $\alpha$-clique extension of a $(\frac{\beta}{\alpha}-[2])\times(r-[2])$-grid. Because of $|\ell\cap M|=\alpha+1$ and $y,z\in\ell\cap M$, it follows that $\ell$ does not contain vertices at distance 3 from $x$. This means that $d(x,\ell)=1$ and hence that $\ell\cap\Gamma(x)\neq\emptyset.$ Conversely, if $\ell$ is a line with $\ell\cap M\neq\emptyset$ and $\ell\cap\Gamma(x)\neq\emptyset$, then $\ell\cap M\subseteq\Gamma_{2}(x)\cap M=B$. This means that $|\ell\cap B|=|\ell\cap M|=\alpha+1\geq 2$. This finishes the proof of the claim.
Let $\ell$ be a line that intersects $B$ in at least two vertices. Then, by Claim 1, we have $|\ell\cap B|=|\ell\cap M|=\alpha+1$.
By Theorem 26, the cardinality of $B$ satisfies $|B|=\phi^{*}_{2}=\alpha(b+1)+1$. For each pair of vertices $y,z\in B$ there exists a unique line $m$ through them. Hence $m\cap B$ is in $L(B)$ and therefore we have shown that $(B,L(B))$ is a 2-$(\alpha(b+1)+1,\alpha+1,1)$-design.
Note that $|L(B)|(\alpha+1)=|B|(b+1)=\alpha(b+1)^{2}+b+1$. It follows that $\alpha+1$ divides $(b+1)^{2}-b-1=b(b+1)$. This shows the theorem.
6 Proof of Theorem 1
In this section we give the proof of Theorem 1.
Proof of Theorem 1. Terwilliger classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ with $D\geq 3$ and $b=1$, and showed that they are the graphs in Items (1)-(4), see [1, Theorem 6.1.1].
Therefore we may assume that $b\geq 2$. If $\alpha=0$, then we are done by Corollary 20. So we assume $\alpha>0$.
We first consider the case that $\alpha\in\{b,b-1\}$. Assume that $\Gamma$ is not a Grassmann graph or a bilinear forms graph. In this case, by [17, Corollary 1.3], we have: if $\alpha=b-1$, then
$\beta$ $<\frac{1}{2b-1}(2b^{4}+2b^{3}+2b^{2}+b-1)r$
$<\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\};$
if $\alpha=b$, then
$\beta$ $<\frac{8}{3}(b^{2}+2b)r$
$<\max\{\frac{2b+4}{2b+3}r(b+2)(\alpha b+b+\alpha),\frac{2b+4}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
From now on, we may assume that
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
By Theorem 19, the graph $\Gamma$ is geometric and hence it has Delsarte clique. This means that $21\leq\alpha\leq b-$ 2 is an integer, as we already have considered $\alpha\in\{0,b-1,b\}$.
Note that $\Gamma$ is geometric. Let $\mathcal{C}$ be a set of Delsarte cliques of $\Gamma$ such that $\Gamma$ is the point graph of the partial linear space $X=(V(\Gamma),{\mathcal{C}},\in)$. By Lemma 22 and Theorem 24 , for each vertex $x$ of $\Gamma$, the local graph $\Delta_{x}$ at vertex $x$ is the $\alpha$-clique extension of a $\frac{\beta}{\alpha}\times r$-grid. By Theorem 28, $\alpha+1$ divides $b(b+1)$.
Acknowledgments
J.H. Koolen is partially supported by the National Natural Science Foundation of China (No. 12471335), and the Anhui Initiative in Quantum Information Technologies (No. AHY150000).