### v2

# Bounding the parameter $\beta$ of a distance-regular graph with classical parameters

Abstract
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfying $b\geq 2$ and $D\geq 3$. Let $r=1+b+b^{2}+\cdots+b^{D-1}$. In 1999, K. Metsch showed that there exists a positive constant $C(\alpha,b)$ only depending on $\alpha$ and $b$, such that if $\beta\geq C(\alpha,b)r^{2}$, then either $\Gamma$ is a Grassmann graph or a bilinear forms graph.
In this work, we show that for $b\geq 2$ and $D\geq 3$, then there exists a constant $C_{1}(\alpha,b)$ only depending on $\alpha$ and $b$, such that if $\beta\geq C_{1}(\alpha,b)r$, then either $\Gamma$ is a Grassmann graph, or a bilinear forms graph.
Key Words: distance-regular graphs, classical parameters, geometric graphs, eigenvalues, cliques
2020 Mathematics Subject Classification. 05E30E-mail addresses: lch1994@mail.ustc.edu.cn (C. Lv), koolen@ustc.edu.cn (J.H. Koolen).
1 Introduction
For undefined notions and more details see [3] and [7]. We introduce distance-regular graphs with classical parameters in the next section. In this paper, we study distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$. It is known that the parameter $b$ is an integer such that $b\neq 0,-1$ (Proposition 7). C. Weng [22] classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\leq-2$ under weak conditions, see Remark 2 below. In this paper we consider the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\geq 1$. The distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b=1$ were classified by Terwilliger, see [3, Theorem 6.1.1]. In [15], Koolen et al. showed that for $b\geq 2$ and $D$ much larger than $b$, the parameter $\alpha$ is bounded by $b(b+1)^{2}+1$. This was improved in [13], where Koolen et al. showed that for $b\geq 2$ and $D\geq 9$, the parameter $\alpha$ is bounded by $b^{2}(b+1)+1$. Moreover, they showed that for $b=2$ and $D\geq 12$, the parameter $\alpha$ is bounded by $2$. Let $r=1+b+\cdots+b^{D-1}$. In [19, Corollary 1.3], Metsch showed that if $b\geq 2$ and $\beta\geq C(\alpha,b)r^{2}$, then the graph must be a bilinear forms graph or a Grassmann graph, where $C(\alpha,b)$ is a constant only depending on $\alpha$ and $b$. Moreover, in the same paper, he showed that, if $\alpha=b-1$, then $\beta<\frac{1}{2b-1}(2b^{4}+2b^{3}+2b^{2}+b-1)r$ or the graph is a bilinear forms graph, and if $\alpha=b$, then $\beta<\frac{8}{3}(b^{2}+2b)r$ or the graph is a Grassmann graph.
In this paper we show that if $b\geq 2$ and $\beta\geq C_{1}(\alpha,b)r$, where $C_{1}(\alpha,b)$ is a constant only depending on $\alpha$ and $b$, then the graph must be a bilinear forms graph, or a Grassmann graph.
Our main result is the following.
Theorem 1
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$, such that $b\geq 1$ and $D\geq 3$. Let $r=1+b+b^{2}+\cdots+b^{D-1}$. Then one of the following holds:
  1. (1)
    $\Gamma$ is a Johnson graph;
  2. (2)
    $\Gamma$ is a Hamming or a Doob graph;
  3. (3)
    $\Gamma$ is a halved cube;
  4. (4)
    $\Gamma$ is the Gosset graph with intersection array $\{27,10,1;1,10,27\}$;
  5. (5)
    $\Gamma$ is a Grassmann graph;
  6. (6)
    $\Gamma$ is a bilinear forms graph;
  7. (7)
    $b\geq 2$ and $\beta<\max\{2\frac{b+2}{2b+3}r(b+2)(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
Remark 1
  1. (1)
    For $\alpha\in\{b-1,b\}$, [19, Corollary 1.3] gives slightly better bounds.
  2. (2)
    For $\alpha=b=2$, the bound of [19, Corollary 1.3] was slightly improved in [14].
  3. (3)
    In [9], they showed that if a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfies $\alpha=b\geq 2$, $D\geq 9$ and $\beta=b(b^{D-1}+b^{D-2}\ldots+1)$, then the graph has to be the Grassmann graph $J_{b}(2D,D)$.
  4. (4)
    In [8], they showed that if a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ satisfies $\alpha+1=b=2$, $D\geq 3$ and $\beta=b^{D}-1$, then the graph has to be the bilinear forms graph $H_{2}(D,D)$.
  5. (5)
    For $b\geq 2$, the twisted Grassmann graphs as found by Van Dam and Koolen in [6] are not geometric and have $\alpha=b$ and $\beta=b^{2}r+b$. This shows that for $\beta=b^{2}r+b$ we can have Item (7) of Theorem 1, and hence the bound in Item (7) is quite tight.
Our proof closely follows the ideas of Metsch as given in [19]. We do not assume that $\alpha$ is an integer at most $b$. Note that for Metsch’s result, it is crucial that $\alpha$ is an integer at most $b$. We obtain this once we show that the distance-regular graph is geometric. The linear algebra idea in the proof of Proposition 19 was first employed in [17].
[…]
2 Definitions and preliminaries
All the graphs considered in this paper are finite, undirected and simple. The reader is referred to [3, 7] for more information. Let $\Gamma$ be a connected graph with vertex set $V(\Gamma)$. The distance $d(x,y)$ between two vertices $x,y\in V(\Gamma)$ is the length of a shortest path between $x$ and $y$ in $\Gamma$. The diameter $D=D(\Gamma)$ of $\Gamma$ is the maximum distance between any two vertices of $\Gamma$. For each $x\in V(\Gamma)$, let $\Gamma_{i}(x)$ be the set of vertices in $\Gamma$ at distance $i$ from $x$ ($0\leq i\leq D$). In addition, define $\Gamma_{-1}(x)=\Gamma_{D+1}(x)=\emptyset$. For the sake of simplicity, we denote $\Gamma_{1}(x)$ by $\Gamma(x)$. For any vertex $x$ of $\Gamma$, the subgraph induced on $\Gamma(x)$ is called the local graph of $\Gamma$ at $x$, and we denote it by $\Delta_{x}$. For a vertex $x$ of $\Gamma$, the cardinality $|\Gamma(x)|$ of $\Gamma(x)$ is called the valency of $x$ in $\Gamma$. In particular, $\Gamma$ is regular with valency $k$ if $k=|\Gamma(x)|$ holds for all $x\in V(\Gamma)$. We also define for vertices $x,y$ of $\Gamma$ at distance $i$ $(i=1,2,\ldots,D)$ the set $C_{i}(x,y)$ by $C_{i}(x,y):=\Gamma_{i-1}(x)\cap\Gamma(y)$. For vertices $u,v$ of $\Gamma$ at distance $j$ $(j=0,1,\ldots,D-1)$, we define the set $B_{j}(u,v)$ by $B_{j}(u,v):=\Gamma_{j+1}(u)\cap\Gamma(v)$.
For a set $T$ of vertices of a connected graph and a vertex $x$, we define the distance $d(x,T)$ by $d(x,T)=\min\{d(x,y)\mid y\in T\}$.
The adjacency matrix $A=A(\Gamma)$ of $\Gamma$ is the matrix whose rows and columns are indexed by vertices of $\Gamma$ and the ($x,y$)-entry is $1$ whenever $x$ and $y$ are adjacent and 0 otherwise. The eigenvalues of $\Gamma$ are the eigenvalues of its adjacency matrix $A$.
A clique of a graph $\Gamma$ is a set of mutually adjacent vertices of $\Gamma$. We sometimes also refer to a complete subgraph of $\Gamma$ as a clique. Let $K_{m,n}$ be the complete bipartite graph with $m$ and $n$ be the number of vertices in each partition.
A connected graph $\Gamma$ is called edge-regular with parameters $(n,k,a)$ if $\Gamma$ has $n$ vertices, is $k$-regular and any two adjacent vertices have exactly $a$ common neighbours. A connected graph $\Gamma$ is called amply-regular with parameters $(n,k,a,c)$ if $\Gamma$ is edge-regular with parameters $(n,k,a)$ and any two vertices at distance 2 have exactly $c$ common neighbours.
2.1 Distance-regular graphs
A graph $\Gamma$ is called distance-regular if there exist integers $b_{i},c_{i}$ $(0\leq i\leq D)$ such that for any two vertices $x,y\in V(\Gamma)$ with $d(x,y)=i$, there are precisely $c_{i}$ neighbors of $y$ in $\Gamma_{i-1}(x)$ and $b_{i}$ neighbors of $y$ in $\Gamma_{i+1}(x)$, where we define $b_{D}=c_{0}=0$. Note that in this case $C_{i}(x,y)$ contains exactly $c_{i}$ vertices for $i=1,2,\ldots,D$ and $B_{i}(x,y)$ contains exactly $b_{i}$ vertices for $i=0,1,\ldots,D-1$. In particular, any distance-regular graph is regular with valency $k:=b_{0}$. We define $a_{i}:=k-b_{i}-c_{i}$ for notational convenience. Note that $a_{i}=|\Gamma(y)\cap\Gamma_{i}(x)|$ holds for any two vertices $x,y$ with $d(x,y)=i$ $(0\leq i\leq D)$ and that the numbers $a_{i},b_{i}$ and $c_{i}$ $(0\leq i\leq D)$ are called the intersection numbers of $\Gamma$. For an eigenvalue $\theta$ of $\Gamma$, the sequence $(u_{i})_{i=0,1,...,D}$ = $(u_{i}(\theta))_{i=0,1,...,D}$ satisfying $u_{0}$ = $u_{0}(\theta)$ = $1$, $u_{1}$ = $u_{1}(\theta)$ = $\theta/k$, and
$c_{i}u_{i-1}+a_{i}u_{i}+b_{i}u_{i+1}=\theta u_{i}~(i=1,2,\ldots,D-1)$ (1)
is called the standard sequence corresponding to the eigenvalue $\theta$ (see [3, p.128]). The following lemma is well-known.
Lemma 2 (cf. [3, Proposition 4.4.1])
Let $\Gamma$ be a distance-regular graph, $\theta\neq k$ an eigenvalue of $\Gamma$, say with multiplicity $m$ and let $(u_{i})_{i}$ be the standard sequence corresponding to $\theta$. Then there exists a map $f:V(\Gamma)\rightarrow\mathbf{R}^{m}$, $x\mapsto\overline{x}$ such that the inner product $(\overline{x},\overline{y})=u_{d(x,y)}$.
The map $f$ is called the standard representation of $\Gamma$ with respect to $\theta$.
2.2 Geometric distance-regular graphs
The following lemma is called the Delsarte bound. We give a proof for it, as we need some consequences later in this paper.
Lemma 3 (cf. [3, Proposition 4.4.6])
Let $\Gamma$ be a distance-regular graph with diameter $D\geq 2$, valency $k$ and smallest eigenvalue $\theta_{\min}$, say with multiplicity $m$. Let $(u_{i})_{i}$ be the standard sequence corresponding to $\theta_{\min}$. Then the order $c$ of a clique $C$ in $\Gamma$ is bounded by $c\leq 1+\frac{k}{-\theta_{\min}}$.
Assume that $c=1+\frac{k}{-\theta_{\min}}$. Then any vertex of $\Gamma$ has distance at most D1D-1 to $C$. Let $x$ be a vertex with $d(x,C)=j$ for some $1\leq j\leq D-1$. Then the number of vertices $\phi_{j}(x)$ in $C$ at distance $j$ from $x$ satisfies:
$(u_{j+1}-u_{j})\phi_{j}(x)=(1+\frac{k}{-\theta_{\min}})u_{j+1},$ (2)
and hence $\phi_{j}(x)$ does not depend on $x$, only on $j$.
Proof. Let $f:V(\Gamma)\rightarrow\mathbf{R}^{m},x\mapsto\overline{x}$ be the standard representation of $\Gamma$ with respect to $\theta_{\min}$. Let $\overline{C}=\sum\overline{u}$ where the sum is taken over all vertices $u$ of $C$. Then the Gram matrix $Gr$ with respect to $\{\overline{u}\mid u$ is a vertex of C}C\} is the matrix that satisfies
$Gr=u_{1}J+(1-u_{1})I.$
As $u_{1}=\frac{\theta_{\min}}{k}$, and $Gr$ has row sum $1+(c-1)u_{1}$ which is non-negative, as $Gr$ is positive semidefinite, we obtain $c\leq 1+\frac{k}{-\theta_{\min}}$, and equality holds if and only if $\overline{C}=\mathbf{0}$. Assume that equality holds. Let $x$ be a vertex of $\Gamma$ such that $d(x,C)=j$. Then the inner product $(\overline{x},\overline{C})$ satisfies
[…]
As $u_{D}\neq 0$, we have $\phi_{D}(x)=0$. This implies that there is no vertex of $\Gamma$ has distance $D$ to $C$. This shows the lemma.
A clique of $\Gamma$ with $1+\frac{k}{-\theta_{\min}}$ vertices is called a Delsarte clique.
A distance-regular graph $\Gamma$ is called geometric if there exists a set of Delsarte cliques $\mathcal{C}$ in $\Gamma$ such that each edge lies in a unique $C\in\mathcal{C}$. In this case, we say that $\Gamma$ is a geometric distance-regular graph with respect to $\mathcal{C}$. Moreover, a partial linear space $X=(V(\Gamma),\mathcal{C},\in)$ (see Section 3 for the definition) is naturally induced, whose point graph is $\Gamma$, and hence the elements of $\mathcal{C}$ are also called lines.
Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$. Let $C$ be a Delsarte clique and $x$ be a vertex at distance $j$ from $C$ for $0\leq j\leq D-1$. Define $\phi_{j}$ be the number of vertices in $C$ at distance $j$ from $x$. For two vertices $x$ and $y$ at distance $j$ for $1\leq j\leq D,$ let $\tau_{j}$ be the number of cliques in $\mathcal{C}$ containing $x$ at distance j1j-1 from $y$. The number $\tau_{j}$ $(j=1,2,,\ldots,D)$ does not depend on the pair $x,y$, only at their distance.
We can express the intersection numbers of $\Gamma$ in terms of the $\phi_{j}$’s and $\tau_{j}$’s as follows (see [12, Lemma 4.1]):
$c_{i}=\tau_{i}\phi_{i-1}\ \ (i=1,2,\ldots,D),\ \ b_{i}=-(\theta_{\min}+\tau_{i})\left(1+\frac{k}{-\theta_{\min}}-\phi_{i}\right)\ \ (i=1,2,\ldots,D-1).$ (3)
We have the following properties for the numbers $(\phi_{i})_{i}$ and $(\tau_{i})_{i}$ for a geometric distance-regular graph.
Lemma 4 (cf. [2, Theorem 5.5])
Let $\Gamma$ be a geometric distance-regular graph with valency $k$ and diameter $D\geq 2$. Suppose $\phi_{1}>1$. Then
$1<\phi_{1}<...<\phi_{D-1}.$
In particular, $D\leq s$, where $s+1$ is the order of a Delsarte clique.
Lemma 5 (cf. [1, Lemma 4.1])
Fix an integer $m\geq 2$. Suppose that $\Gamma$ is a geometric distance-regular graph with diameter $D\geq 2$ and smallest eigenvalue m-m. If $\phi_{1}\geq 2$, then
$2\leq\phi_{1}\leq\tau_{2}<...<\tau_{D}=m.$
Lemma 6 (cf. [2, Lemma 5.2])
Let $\Gamma$ be a distance-regular graph with valency $k$ and diameter $D\geq 2$, containing a Delsarte clique $C$ of size $s+1$. Then the following hold:
  1. (1)
    Let $i$ and $j$ be positive integers such that $i+j\leq D-1$. Then $\phi_{i}+\phi_{j}\leq s+1$.
  2. (2)
    If $\phi_{1}>\frac{s+1}{2}$, then $D=2$.
2.3 Distance-regular graphs with classical parameters
A distance-regular graph $\Gamma$ is said to have classical parameters $(D,b,\alpha,\beta)$ if the diameter of $\Gamma$ is $D$ and the intersection numbers of $\Gamma$ can be expressed as follows:
$b_{i}=([D]-[i])(\beta-\alpha[i]),~~0\leq i\leq D-1,$
$c_{i}=[i](1+\alpha[i-1]),~~1\leq i\leq D,$
where
[…]
From [3, Corollary 8.4.2], we know that the eigenvalues of $\Gamma$ are
$D-i](\beta-\alpha[i])-[i]=\frac{b_{i}}{b^{i}}-[i],~~0\leq i\leq D.$
We note that $c_{2}=(b+1)(\alpha+1)$ and that if $b\geq 1$, then the eigenvalues $\theta_{i}=\frac{b_{i}}{b^{i}}-[i](0\leq i\leq D)$ of $\Gamma$ are in the natural ordering, i.e., $k=\theta_{0}>\theta_{1}>\cdots>\theta_{D}$.
From the following result we know that the parameter $b$ of a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ is an integer.
Proposition 7 (cf. [3, Proposition 6.2.1])
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ and diameter $D\geq 3$. Then $b$ is an integer such that $b\neq 0,-1$.
The following lemma is an easy consequence of Proposition 7.
Lemma 8
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ and diameter $D\geq 3$, such that $b\geq 1$. Then $\alpha\geq 0$.
Proof. As $b$ is a positive integer and $c_{2}=(\alpha+1)(b+1)$ is clearly a positive integer, we have $\alpha(b+1)$ is an integer. Now the intersection number $c_{3}$ satisfies $c_{3}=(1+b+b^{2})(1+(1+b)\alpha)\geq 1$. This means $(1+b)\alpha+1>0$, and as it is an integer, we see $(1+b)\alpha+1\geq 1$. This shows the lemma.
The members of many of the infinite families of the known distance-regular graphs with unbounded diameter have classical parameters, like the Johnson graphs, Hamming graphs, Grassmann graphs and the bilinear forms graphs where the parameter $b$ is positive.
Remark 2
  1. (1)
    For $b\leq-2$, C. Weng [22] classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ and $b\leq-2$. His result states: If $D\geq 4$, $a_{1}\neq 0$ and $c_{2}>1$, then one of the following holds: $(1)$ $\Gamma$ is the dual polar graph $^{2}A_{2D-1}(-b)$, $(2)$ $\Gamma$ is the Hermitian forms graph $Her_{-b}(D)$, $(3)$ $\alpha=\frac{b-1}{2}$, $\beta=-\frac{1+b^{D}}{2}$ and b-b is a power of an odd prime.
  2. (2)
    Tian et al. [21] give the current state of art on distance-regular graphs with classical parameters with $b\leq-2$.
Before we look at the Delsarte bound for distance-regular graphs with classical parameters, we will determine the standard sequence for the smallest eigenvalue of such a distance-regular graph. Let $b\geq 2$ be an integer and let $i]=\frac{b^{i}-1}{b-1}$ and $r=[D]$, as before.
Lemma 9
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $b\geq 2$ and $D\geq 2$. The standard sequence $(u_{i})_{i}$ corresponding to the smallest eigenvalue r-r of $\Gamma$ satisfies: $u_{0}=1$ and $u_{i}=\frac{(-1)^{i}}{\beta}\frac{(1+\alpha[1])(1+\alpha[2])\cdots(1+\alpha[i-1])}{(\beta-\alpha[1])(\beta-\alpha[2])\cdots(\beta-\alpha[i-1])}$ for $i\geq 1$.
Proof. We have $c_{i}=[i](1+\alpha[i-1])$ for $1\leq i\leq D$, $a_{i}=[i](\beta-1+\alpha(r-[i]-[i-1]))$ for $1\leq i\leq D$ and $b_{i}=(r-[i])(\beta-\alpha[i])$ for $0\leq i\leq D-1$. Note that $k=b_{0}=\beta r$. The $u_{i}$’s satisfy $u_{0}=1$, $u_{1}=\frac{-r}{k}=\frac{-1}{\beta}$ and $c_{i}u_{i-1}+a_{i}u_{i}+b_{i}u_{i+1}=-ru_{i}$ for $1\leq i\leq D-1$. The lemma follows by an easy induction.
The following result is a direct consequence of Lemmas 3 and 9 for a distance-regular graph with classical parameters.
Lemma 10
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ with $D\geq 3$ and $b\geq 2$. Then the order $c$ of a clique $C$ in $\Gamma$ is bounded by $c\leq\beta+1$. If equality holds, then the number $\phi_{j}$ for a vertex at distance $j$ from $C$ satisfies $\phi_{j}=1+\alpha[j]$ for $0\leq j\leq D-1$. In particular $\phi_{1}=1+\alpha$.
Proof. As the smallest eigenvalue $\theta_{\min}$ of $\Gamma$ is equal to r-r where $r=[D]$, and $k=\beta r$ we obtain that any clique $C$ with order $c$ satisfies $c\leq 1+\frac{k}{r}=1+\beta$. If equality holds then by Equation (2) we obtain $\phi_{j}(x)=1+\alpha[j]$, as the standard sequence corresponding to eigenvalue r-r satisfies $u_{i}=\frac{(-1)^{i}}{\beta}\frac{(1+\alpha[1])(1+\alpha[2])\cdots(1+\alpha[i-1])}{(\beta-\alpha[1])(\beta-\alpha[2])\cdots(\beta-\alpha[i-1])}$ for $i\geq 1$, by Lemma 9. This shows the lemma.
Now we can calculate the $\phi_{j}$’s and $\tau_{j}$’s of Equation (3) for a geometric distance-regular graph with classical parameters.
Lemma 11
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$, and assume that $\Gamma$ is geometric. Then the numbers $\phi_{j}$’s and $\tau_{j}$’s of Equation (3) satisfy
$\phi_{j}=1+\alpha[j],\ \ (j=0,1,\ldots,D-1),\ \ \tau_{j}=[j]\ \ (j=1,2,\ldots,D),$
where $j]=\frac{b^{j}-1}{b-1}$.
Proof. The expressions for the $\phi_{j}$’s follow from Lemma 10 and for the $\tau_{j}$’s follow now from Equation (3).
Now we first consider $\alpha=0$.
Proposition 12
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 1$, $\alpha=0$, $\beta>1$ and $D\geq 3$. If $\Gamma$ is geometric, then $\Gamma$ is a dual polar graph or a Hamming graph.
Proof. As $\alpha=0$ we have $a_{1}=\beta-1>0$ and $\Gamma$ is locally the disjoint union of cliques of order $\beta$. By [3, Theorem 9.4.4], we are now done.
2.4 Strongly regular graphs
A graph $\Gamma$ is called a strongly regular graph with parameters $(n,k,\lambda,\mu)$ if $\Gamma$ has $n\geq 2$ vertices, is $k$-regular and any two distinct vertices have $\lambda$ (resp. $\mu$) common neighbours depending on whether they are adjacent or not. For more information and for proofs of some the results we mention without proof see [11, Chapter 10] and [4].
If $\Gamma$ or its complement is disconnected, we say $\Gamma$ is imprimitive, and primitive otherwise. Note that the only connected imprimitive non-complete strongly regular graphs are the complete multipartite graphs $K_{m\times s}$ with $m\geq 2,s\geq 2$. It is known that a connected non-complete $k$-regular graph is strongly regular if and only if it has exactly three distinct eigenvalues $k>\sigma_{1}>\sigma_{2}$. Note that a non-complete connected strongly regular graph with parameters $(n,k,\lambda,\mu)$ has non-integral eigenvalues only if $n=4\mu+1,k=2\mu,$ and $\lambda=\mu-1.$
We say that a connected non-complete strongly regular graph with parameters $(n,k,\lambda,\mu)$ has classical parameters $(b,\alpha,\beta)$ if $k=\beta(b+1)$, $\lambda=\beta-1+\alpha b$ and $\mu=(b+1)(\alpha+1)$. Note that a distance-regular graph with classical parameters $(2,b,\alpha,\beta)$ is just a strongly regular graph with classical parameters $(b,\alpha,\beta)$. Also note that any connected non-complete strongly regular graph with distinct eigenvalues $k>\sigma_{1}>\sigma_{2}$ has classical parameters $(b,\alpha,\beta)$ where $b+1\in\{-\sigma_{1},-\sigma_{2}\}$.
We now reformulate Theorem 8.6.3 of [4] in terms of terms of $(b,\alpha,\beta)$. Note that, if $b>0$, then the smallest eigenvalue is equal to b1-b-1. We refer to [4, Sections 8.4.2, 8.6] for the definition of the block graph of a 2-$(v,k,1)$-design and the Latin square graph $LS_{m}(n)$.
Theorem 13
Let $b$ be an integer at least 2. Let $\Gamma$ be a primitive strongly regular graph with classical parameters $(b,\alpha,\beta)$. Let $f(b,\alpha)=\frac{1}{2}b(b+1)((\alpha+1)(b+1)+1)+\alpha$. Then $\beta\leq f(b,\alpha)$ or one of the following hold:
  1. (1)
    $\alpha=b$ and $\Gamma$ is the block graph of a 2-$(b\beta+b+1,b+1,1)$-design;
  2. (2)
    $\alpha=b-1$ and $\Gamma$ is a Latin square graph $LS_{b+1}(\beta+1)$.
3 Partial linear spaces
An incidence structure is a tuple $(\mathcal{P},\mathcal{L},\mathcal{I})$ where $\mathcal{P}$ and $\mathcal{L}$ are non-empty disjoint sets and $\mathcal{I}\subseteq\mathcal{P}\times\mathcal{L}$. The elements of $\mathcal{P}$ and $\mathcal{L}$ are called points and lines, respectively. If $(p,\ell)\in\mathcal{I}$, then $p$ is said to be incident with $\ell$. In this case, we also say that $\ell$ contains $p$ or that $p$ lies on $\ell$. The order of a point is the number of lines it is incident with and similarly for lines. The point-line incidence matrix of $(\mathcal{P},\mathcal{L},\mathcal{I})$ is the $|\mathcal{P}|\times|\mathcal{L}|$-matrix such that the $(p,\ell)$-entry is 1 if $p$ is incident with $\ell$ and 0 otherwise. If it is clear from the context what the incidence relation is, then we omit it.
A partial linear space is an incidence structure such that each pair of distinct points are both incident with at most one line.
Let $X=(\mathcal{P},\mathcal{L},\mathcal{I})$ be a partial linear space. For a non-incident pair $(x,\ell)$ of a point $x$ and line $\ell$, let $\tau(x,\ell)$ be the number of points on $\ell$ that are collinear with $x$. We define $\tau(X)$ for a partial linear space $X=(\mathcal{P},\mathcal{L},\mathcal{I})$, as the number $\tau(X)=\max\{\tau(x,\ell)\mid x\in\mathcal{P},\ell\in\mathcal{L},\mbox{ and }(x,\ell)\notin\mathcal{I}\}$.
The point graph $\Gamma$ of an incidence structure $(\mathcal{P},\mathcal{L},\mathcal{I})$ is the graph with vertex set $\mathcal{P}$ and two distinct points are adjacent if and only if they are on a common line. Given a line $\ell$, the vertex set $\{x\in\mathcal{P}\mid(x,\ell)\in\mathcal{I}\}$ induces cliques in $\Gamma$. If the point graph is connected, then for a vertex $x$ and a line $\ell$, we define the distance $d(x,\ell)$ by $d(x,\ell)=\min\{d(x,y)\mid y\in\ell\}$. For two vertices $x,y$ at distance 2 in $\Gamma$, we denote by $[x,y]$ the set of lines $\ell$ through $x$ such that $d(y,u)\leq 2$ for all $u\in\ell$. By Lemma 3, the set of lines $[x,y]$ are exactly the lines through $x$ at distance 1 from $y$ for a geometric distance-regular graph.
First we give the so-called claw bound. It was first proven by Metsch [18, Lemma 1.1] and was rediscovered by Godsil [10, Lemma 2.3] and Koolen and Park [16, Lemma 2].
Lemma 14
Let $\Gamma$ be an amply-regular graph with parameters $(n,k,\lambda,\mu)$. Assume that there exists a positive integer $s$ such that the following holds:
$(s+1)(\lambda+1)-k>(\mu-1)\binom{s+1}{2}$.
Then $\Gamma$ does not contain an induced $K_{1,s+1}$.
Now we recall the following result of Metsch. Note that the first condition is just the claw bound.
Theorem 15 (cf. [19, Result 2.1])
Let $\Gamma$ be an amply-regular graph with parameters $(n,k,\lambda,\mu)$. Assume that there exists a positive integer $s$ such that the following two conditions are satisfied:
  1. (1)
    $(s+1)(\lambda+1)-k>(\mu-1)\binom{s+1}{2}$;
  2. (2)
    $\lambda+1>(\mu-1)(2s-1)$.
Define a line as a maximal clique with at least $\lambda+2-(\mu-1)(s-1)$ vertices. Then $X=(V(\Gamma),{\mathcal{L}},\in)$ is a partial linear space, where $\mathcal{L}$ is the set of all lines, and $\Gamma$ is the point graph of $X$. Moreover, each vertex lies on at most $s$ lines.
4 The SPLS(s)(s) Property
[…]
We have $\gamma\leq 1+\frac{k}{-\theta_{\min}}$ where $\theta_{\min}$ is the smallest eigenvalue of $\Gamma$, and equality if and only if $\Gamma$ is geometric.
Let $\Gamma$ be a distance-regular graph that is the point graph of a partial linear space $X=(V(\Gamma),{\mathcal{L}},\in)$. A vertex $x$ of $\Gamma$ is called Delsarte vertex if all the lines in $\mathcal{L}$} through $x$ are Delsarte cliques.
Lemma 16
[…]
Proof. By Lemma 10, the order $c$ of a clique $C$ in $\Gamma$ satisfies $c\leq 1+\beta$ with equality if and only if $C$ is a Delsarte clique. Now $k=\beta r$. By the definition of a partial linear space the lemma follows.
Now we prepare a lemma that will bound the parameter $\alpha$ of a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$.
Lemma 17
[…]
Proof. Let $x,z$ be two vertices at distance two from each other such that $x$ is a Delsarte vertex and there exists a Delsarte clique $C^{\prime}$ containing $z$ such that $x$ has at least one neighbour in $C^{\prime}$ and hence $x$ has exactly $\alpha+1$ neighbours in $C^{\prime}$. This means that $\alpha+1$ is a positive integer. Let $r=[D]$. Let $C_{1},C_{2},\ldots,C_{r}$ be the lines through $x$, all of which are Delsarte cliques.
The vertices $x$ and $z$ have $c_{2}=(1+\alpha)(1+b)$ common neighbours and $z$ has exactly $\alpha+1$ or 0 neighbours in each of the $C_{i}$ for $i=1,2,\ldots,r$. This means that there are exactly $b+1$ $C_{i}$’s such that $z$ has $\alpha+1$ neighbours in $C_{i}$. Without loss of generality we may assume that $z$ has $\alpha+1$ neighbours in $C_{i}$ for $i=1,2,\ldots,b+1$. On the other hand $x$ has exactly $\alpha+1$ neighbours in $C^{\prime}$. Now the clique $C_{i}$ and $C^{\prime}$ intersect in at most one vertex for $i=1,2,\ldots,b+1$ and they do not intersect if $i\geq b+2$. This means that $\alpha+1\leq b+1$, and hence $\alpha\leq b$.
Let $y$ be a Delsarte vertex in $\Gamma$ at distance at most two from $x$ and distinct from $x$.
Assume first that $x\sim y$. There exists a Delsarte clique $C^{\prime}$ that contains $y$ but not $x$. As $C^{\prime}$ is a maximal clique there exists a vertex $z$ in $C^{\prime}$ not adjacent to $x$. So then we obtain $\alpha\leq b$ by the above.
Now assume $d(x,y)=2$. Then we take $z=y$, and we are done again.
[…]
We say that a distance-regular graph $\Gamma$ has the SPLS(c,s)(c,s) property for some integers $c\geq 1$ and $s\geq 2$, if $\Gamma$ is the point graph of a partial linear space $X=(V(\Gamma),{\mathcal{L}},\in)$, where the set of lines $\mathcal{L}$ are the maximal cliques with at least $a_{1}+2-(c-1)(s-1)$ vertices, $a_{1}\geq(2s-1)(c-1)$, each vertex lies on at most $s$ lines and $\tau(X)\leq c$.
[…]
Note that if a distance-regular graph $\Gamma$ satisfies the conditions of Theorem 15 for a positive integer $s$, then it has the SPLS(s)(s) property. Note that the twisted Grassmann graphs $\tilde{J}_{q}(2D+1,D)$ have the PLS($\frac{q^{D+1}-1}{q-1})$ property (see [6]) , but do not have the SPLS(s)(s) property for any $s\geq 2$. This follows from the Proposition 19, as the twisted Grassmann graphs do not have Delsarte vertices.
Now we give some bounds for $s$ for a distance-regular graph satisfying the SPLS(s)(s) property.
Lemma 18
[…]
Proof. As each clique has at most $1+\frac{k}{-\theta_{\min}}$ vertices, it is clear that $\theta_{\min}\leq s$ and if equality holds, then each line is a Delsarte clique and hence the graph is geometric.
Let $x$ be a vertex that lies on exactly $\sigma$ lines. Then it is easy to show that $x$ lies in an induced $K_{1,\sigma}$, by an easy induction argument as any line has at least $(c_{2}-1)\sigma+2$ vertices and that $\Gamma$ does not contain an induced $K_{1,\sigma+1}$. This shows the lemma, by applying Lemma 14.
Now we give an extra condition on distance-regular graphs with classical parameters that have the SPLS(s)(s) property.
Proposition 19
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $s$ be a positive integer and assume $\Gamma$ has the SPLS(s)(s) property. Let $r=[D]$. Then the following hold:
  1. (1)
    $s\geq r$;
  2. (2)
    more than $\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$ vertices are Delsarte vertices;
  3. (3)
    $\alpha$ is an integer such that $0\leq\alpha\leq b$ and $\beta$ is a positive integer;
  4. (4)
    There exists two vertices $x,y$ at distance 2 such that $|[x,y]|\geq b+1$.
Proof. The smallest eigenvalue of $\Gamma$ is equal to r-r, and therefore $s\geq r$. By the SPLS(s)(s) property, let $X=(V(\Gamma),\mathcal{C},\in)$ be the partial linear space with $\Gamma$ as its point graph, where $\mathcal{C}$ is the set of maximal cliques that have at least $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2$ vertices. Note that all the Delsarte cliques of $\Gamma$ are in $\mathcal{C}$, as all cliques in $\Gamma$ have at most $\beta+1$ vertices and clearly $a_{1}+2-(s-1)(c_{2}-1)\leq\beta+1$, as otherwise $\mathcal{C}$ would be empty. Let $N$ be the point-line incidence matrix of $X$. Let $A$ be the adjacency matrix of $\Gamma$. Now $NN^{T}=A+Q$ where $Q$ is a diagonal matrix. Note that the rank of $N$ equals the rank of $A+Q$.
By Lemma 16 we see that a vertex $x$ of $\Gamma$ lies on exactly $r$ lines if and only if $x$ is a Delsarte vertex. This means that for all vertices $x$ of $\Gamma$ the entry $Q_{xx}\geq r$ and $Q_{xx}=r$ if and only if $x$ is a Delsarte vertex.
Let $\delta$ be the dimension of the null-space of $A+Q$ and $\kappa$ be the number of Delsarte vertices. Let $\mathbf{u}\neq\mathbf{0}$ such that $(A+Q)\mathbf{u}=\mathbf{0}$. We find $\mathbf{u}^{T}(A+rI)\mathbf{u}\geq 0$, as $A$ has smallest eigenvalue r-r, and $\mathbf{u}^{T}(Q-rI)\mathbf{u}\geq\mathbf{0}$, as $Q_{xx}\geq r$ for all vertices $x$ of $\Gamma$. This implies that $A\mathbf{u}=-r\mathbf{u}$ and $Q\mathbf{u}=r\mathbf{u}$. Hence we obtain that the multiplicity $m$ of $r$ as an eigenvalue of $Q$ is at least $\delta$. Further we have that $m$ is exactly the number of vertices $x$ such that $Q_{xx}=r$. This shows that $\delta\leq\kappa$.
Now we determine the rank of $N$. This is at most the number of lines. So we need to estimate the number $\eta$ of lines. We do this by double counting the pairs $(x,\ell)$ where $x$ is a vertex of $\Gamma$ and $\ell$ is a line. Let $w:=|\{(x,\ell)\mid x\in V(\Gamma),\ell\in\mathcal{C},x\in\ell\}|.$ On the one hand we have that $w\leq s|V(\Gamma)|$ as every vertex lies on at most $s$ lines. On the other hand we find $w\geq\eta(a_{1}+2-(s-1)(c_{2}-1))$ as a line contains at least $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2$ vertices. Hence we obtain that
$\eta\leq\frac{s|V(\Gamma)|}{a_{1}+2-(s-1)(c_{2}-1)}.$
As $a_{1}+2-(s-1)(c_{2}-1)\geq s(c_{2}-1)+2>s(c_{2}-1),$ we find that $\eta<\frac{1}{c_{2}-1}|V(\Gamma)|.$ So the rank of $N$ is less than $\frac{1}{c_{2}-1}|V(\Gamma)|$. This means that the dimension $\delta$ of the null-space of $A+Q$ is more than $|V(\Gamma)|-\frac{1}{c_{2}-1}|V(\Gamma)|=\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$. As $b\geq 2$ and $\alpha\geq 0$, by Lemma 8, we have $c_{2}=(\alpha+1)(b+1)\geq 3$. Then, as the number $\kappa$ of Delsarte vertices is at least the dimension $\delta$ and this number is in turn more than $|V(\Gamma)|-\frac{1}{c_{2}-1}|V(\Gamma)|=\frac{c_{2}-2}{c_{2}-1}|V(\Gamma)|$, we find more than half the vertices are Delsarte vertices. It means that there are adjacent vertices which are both Delsarte vertices. By Lemma 17 we find that $\alpha\geq 0$ is an integer and $\alpha\leq b$. If we take $x$ as a Delsarte vertex and $y$ a vertex at distance 2 of $x$, then $y$ has $\alpha+1$ or 0 neighbours on any line through $x$. As $c_{2}=(\alpha+1)(b+1)$, we see that there are at least $b+1$ lines in $[x,y]$. This shows the proposition.
Now we give a sufficient condition for a distance-regular graph with classical parameters to have the SPLS(s)(s) property.
Theorem 20
[…]
$\beta>\frac{s+1}{2(s+1-r)}\left((\alpha b+\alpha+b)s-2\alpha(r-1)\right),$ (4)
and
$\beta>(\alpha b+b+\alpha)2s-\alpha r-(\alpha+1)b.$ (5)
Let $s_{0}:=\frac{4r}{3}$ be a rational number. Then
$\frac{s+1}{s+1-r}<\frac{s_{0}}{s_{0}-r}=4.$
So to show that Conditions (4) and (5) hold, it suffices to show that the following two conditions hold:
$\beta>2(\alpha b+b+\alpha)s_{0}-4\alpha(r-1)=\frac{4}{3}(2(\alpha b+b+\alpha)r-3\alpha(r-1))$ (6)
and
$\beta>2(\alpha b+b+\alpha)s_{0}-\alpha r-(\alpha+1)b=\frac{1}{3}(8(\alpha b+b+\alpha)r-3\alpha r-3(\alpha+1)b).$ (7)
When $\beta\geq\frac{1}{3}(8\alpha b+8b+5\alpha)r$, Conditions (6) and (7) clearly hold. This shows the theorem.
5 Geometric distance-regular graphs
In this section we will give a sufficient condition for a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ which has the SPLS(s)(s) property to be geometric.
Our main result in this section is the following and we will give the proof using several claims.
Theorem 21
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $s$ be a positive integer and assume $\Gamma$ has the SPLS(s)(s) property. Let $r=[D]$. Assume that $\beta$ satisfies the following bounds:
  1. (1)
    $\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha$;
  2. (2)
    $\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$;
  3. (3)
    $\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b$.
Then $\Gamma$ is geometric.
Proof. Note that by Proposition 19, we have $s\geq r$ and $0\leq\alpha\leq b$ where $\alpha$ is an integer. Note that by (1), we have $\beta>(b^{2}+b+1)(b+1)$ as $0\leq\alpha\leq b$ and $D\geq 3$.
Now we will look at the line set $[x,y]$ for two vertices $x,y$ at distance 2. First, we show that this set has cardinality at most $2b+1$, and later, we will improve it to $b+1$.
[…]
Proof of Claim 1.
Let $M$ be the set of vertices consisting of neighbours of $x$ at distance at most 2 from $y$. Then
$|M|$ $=k-b_{2}=\beta r-\left(r-\frac{b^{2}-1}{b-1}\right)\left(\beta-\alpha\frac{b^{2}-1}{b-1}\right)$
$=(b+1)(\beta+r\alpha-(b+1)\alpha)=(b+1)(a_{1}+1-b\alpha).$
As every line has at least $a_{1}+2-(s-1)(c_{2}-1)$ vertices, every line of $[x,y]$ has at least $a_{1}+1-(s-1)(c_{2}-1)$ vertices in $M$, because $x\not\in M$.
Therefore
$|[x,y]|$ $\leq\frac{|M|}{a_{1}+1-(s-1)(c_{2}-1)}$ (8)
$=\frac{(b+1)(a_{1}+1-b\alpha)}{a_{1}+1-(s-1)(c_{2}-1)}=(b+1)\left(1+\frac{(s-1)(c_{2}-1)-b\alpha}{a_{1}+1-(s-1)(c_{2}-1)}\right).$ (9)
If $\alpha\geq 1$, by the SPLS(s) property we have $a_{1}+1\geq(2s-1)(c_{2}-1)+1=2(s-1)(c_{2}-1)+c_{2}$, and hence we have $|[x,y]|<2(b+1)$. If $\alpha=0$, then, by Condition (1) of the theorem, we have
$|[x,y]|\leq\frac{(b+1)\beta}{\beta-(s-1)b}<b+2.$
This shows the claim.
Claim 2
  1. (1)
    We have $|[x,y]|\leq b+1$ for any two vertices $x$ and $y$ of $\Gamma$ at distance 2;
  2. (2)
    If a vertex has distance 1 to a line, then it has at most $b+1$ neighbours on the line;
  3. (3)
    $\Gamma$ has the SPLS(b+1,s)(b+1,s) property.
Proof of Claim 2. By Claim 1, we may assume $\alpha\geq 1$. Let $q$ be the maximum integer such that there exist vertices $x,y$ at distance 2 with $|[x,y]|=q+1$. By Proposition 19 we have $q\geq b$.
We will first show that every line has at least $a_{1}+2-(s-1)q$ vertices. Let $\ell$ be a line and let $u\sim v$ be vertices on $\ell$. We’ll show that any line $\ell_{1}$ through $u$ different from $\ell$ contains at most $q$ neighbours of $v$ other than $u$. Let $w$ on $\ell_{1}$ at distance 2 from $v$. Each neighbour of $v$ other than $u$ on the line $\ell_{1}$ gives a unique line in $[v,w]$. This means that $\ell_{1}$ contains at most $q$ neighbours of $v$ besides $u$. As $u$ lies on at most $s$ lines, we obtain that $u$ and $v$ have at most $|\ell|-2+(s-1)q$ common neighbours. Hence $a_{1}\leq|\ell|-2+(s-1)q$, or equivalently, $|\ell|\geq a_{1}+2-(s-1)q$. Let $M$ be as in Claim 1. As every line has at least $a_{1}+2-(s-1)q$ vertices, every line of $[x,y]$ has at least $a_{1}+1-(s-1)q$ vertices in $M$. Therefore $(q+1)(a_{1}+1-(s-1)q)\leq|M|$. Let $f(x)=(x+1)(a_{1}+1-(s-1)x)$. Then $f(q)\leq|M|$. By Claim 1 we have $q\leq 2b$, and we have already seen $q\geq b$. We will show that $q=b$. Note that $|M|=(b+1)(a_{1}+1-b\alpha).$ The minimum of $f(x)$ for $b+1\leq x\leq 2b$ is the minimum of $f(b+1)$ and $f(2b)$.
First we show that $f(b+1)\leq f(2b)$ holds.
Now
$f(2b)-f(b+1)=(2b+1)(a_{1}+1-(s-1)2b)-(b+2)(a_{1}+1-(s-1)(b+1)).$
This implies that
$f(2b)-f(b+1)=(b-1)(a_{1}+1)-(s-1)(3b^{2}-b-2)=(b-1)(a_{1}+1-(s-1)(3b+2)).$
So $f(2b)\geq f(b+1)$ if $a_{1}+1\geq(s-1)(3b+2)$. We already found that
$a_{1}+1\geq 2(s-1)(c_{2}-1)+c_{2}=2(s-1)((b+1)(\alpha+1)-1)+(b+1)(\alpha+1).$
We see that $f(2b)\geq f(b+1)$ as $\alpha\geq 1$ and $b\geq 2$. Now we need to compare $f(b+1)$ with $|M|$. We find
$f(b+1)-|M|$ $=(b+2)(a_{1}+1-(s-1)(b+1))-(b+1)(a_{1}+1-b\alpha)$
$=a_{1}+1-(b+1)((b+2)(s-1)-b\alpha).$
As $a_{1}+1=\beta+\alpha(r-1)$ and $\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha$, we have
$f(b+1)>|M|.$
This means that $q\leq b$ and hence this shows Item (1) of this claim.
Let $x$ be a vertex and $\ell$ a line such that $d(x,\ell)=1$. Then $x$ has at most $c_{2}$ neighbours on $\ell$. This means that there exists a vertex $y$ on $\ell$ at distance 2 from $x$. There are at most $b+1$ lines in $[x,y]$ and each of them intersects in at most one vertex with $\ell$ and a line through $x$ that intersects with $\ell$ is in $[x,y]$. So $x$ has at most $b+1$ neighbours in $\ell$.
Now it follows that $\Gamma$ has the SPLS(b+1,s)(b+1,s) property, as through each vertex there at most $s$ lines. This shows the claim.
Now we will look at vertices at distance 3 in $\Gamma$.
Claim 3 Let $x,y$ be two vertices at distance 3 in $\Gamma$, and $\ell$ a line through $y$. Then the number of vertices $z$ at distance 2 from $x$ such that $z\in\ell$ is less than $(b+1)(b^{2}+b+1)$.
Proof of Claim 3. Let $U=\Gamma_{2}(x)\cap\ell$ and $W=\Gamma(x)\cap\Gamma_{2}(y)$. Then $|W|=c_{3}=(b^{2}+b+1)(1+\alpha(b+1)$. If $u\in U$, then $\Gamma(u)\cap\Gamma(x)\subseteq W$. Hence every vertex of $U$ has $c_{2}$ neighbours in $W$. Counting the number of edges between $U$ and $W$ we find $|U|c_{2}=\sum_{w\in W}|\Gamma(w)\cap\ell|.$ As a vertex outside $\ell$ has at most $b+1$ neighbours on $\ell$, we obtain $|U|c_{2}\leq|W|(b+1)=(b+1)(b^{2}+b+1)(1+\alpha(b+1))$. Now as $c_{2}=(b+1)(\alpha+1)$, we see that $|U|\leq\frac{(b+1)(b^{2}+b+1)(1+\alpha(b+1))}{(b+1)(\alpha+1)}<(b^{2}+b+1)(b+1)$, and this shows the claim.
Claim 4 If $\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$, then for any pair of vertices $x,y$ at distance 2, the number $|[x,y]|$ satisfies $|[x,y]|=b+1$.
Proof of Claim 4. As above let $M$ be the set of neighbours of $x$ that have distance 1 or 2 from $y$. Then $|M|=k-b_{2}=(b+1)\beta+\alpha(b+1)r-\alpha(b+1)^{2}$. Let $c:=(b+1)(b^{2}+b+1).$ As $\Gamma$ has the SPLS(b+1,s)(b+1,s) propert, by Claim 2, we only need to show that $|[x,y]|\geq b+1$. Assume that $|[x,y]|\leq b$. Any line through $x$ contains at most $\beta+1$ vertices, and therefore, there are at most $\beta$ vertices in $M$.
A line through $x$ that is not in $[x,y]$ contains a vertex at distance 3 from $y$, and hence by Claim 3 contains at most c2c-2 vertices in $M$. As $x$ is on at most $s$ lines, it follows that $|M|\leq|[x,y]|\beta+(s-|[x,y]|)(c-2)\leq b\beta+(s-b)(c-2)$, as $c-2<\beta$. As $|M|=k-b_{2}=(b+1)\beta+\alpha(b+1)r-\alpha(b+1)^{2}$, we obtain $\beta\leq(s-b)(c-2)-\alpha(b+1)r+\alpha(b+1)^{2}$. So, if $\beta>(s-b)(c-2)-\alpha(b+1)r+\alpha(b+1)^{2}=((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2}$, then $|[x,y]|=b+1$, as we have $s\geq r\geq b^{2}+b+1$ and $\alpha\leq b$. This shows Claim 4.
Claim 5 If $\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b$, then for each pair of a vertex $x$ and a line $\ell$, with $d(x,\ell)=1$ and vertices $y_{1},y_{2}$ on $\ell$ with $d(x,y_{1})=d(x,y_{2})=2$, we have that the sets of lines $x,y_{1}]$ and $x,y_{2}]$ are equal.
Proof of Claim 5. Assume that there are vertices $y_{1},y_{2}\in\Gamma_{2}(x)$ that are both on $\ell$ such that $x,y_{1}]\neq[x,y_{2}]$. Put $S_{i}:=\Gamma_{i}(x)\cap\ell$ and $n_{i}:=|S_{i}|$ for $i=1,2.$ Further, put $T:=[x,y_{1}]\cap[x,y_{2}]$ and $t:=|T|$. Then $1\leq n_{1}\leq t\leq b$, as two distinct lines intersect in at most one vertex, $x,y_{1}]\neq[x,y_{2}]$ and $|[x,y_{1}]|=|[x,y_{2}]|=b+1$. Let $T_{2}$ be the set of lines through $x$ that are not in $T$. Then $t_{2}:=|T_{2}|\leq s-t$, since $x$ lies on at most $s$ lines. Let $\ell_{2}\in T_{2}$ be a line through $x$ that is not in $x,y_{1}]$. Then there exists a vertex $u$ on $\ell_{2}$ at distance 3 from $y_{1}$. Now we obtain $|\Gamma_{2}(u)\cap\ell|\leq c-1$ where $c=(b+1)(b^{2}+b+1)$. As $u\sim x$ and $d(u,y_{1})=3$, we have $S_{1}\subseteq\Gamma_{2}(u)\cap\ell$. As a consequence we find $|\Gamma_{2}(u)\cap S_{2}|\leq|\Gamma_{2}(u)\cap\ell|-|S_{1}|\leq c-1-n_{1}.$ It follows that there are at most $c-1-n_{1}$ vertices $v\in S_{2}$ with $\ell_{2}\in[x,v]$, as $d(u,v)\leq 2$ for such a vertex $v$.
On the other hand, if $v\in S_{2}$, then $[x,v]$ contains at least b+1tb+1-t lines of $T_{2}$. By counting pairs $(v,\ell_{2})$ with $v\in S_{2}$ and $\ell_{2}\in T_{2}\cap[x,v]$ we obtain
$n_{2}(b+1-t)\leq|T_{2}|(c-1-n_{1})\leq(s-t)(c-1-n_{1}).$
As $t\leq b$ and $s\geq r>b+1$, we obtain $(b-t)(s-b-1)\geq 0$ which is equivalent with $s-t\leq(b+1-t)(s-b)$. As a consequence we have $n_{2}\leq(s-b)(c-1-n_{1})$. This means that $|\ell|=n_{1}+n_{2}\leq n_{1}+(s-b)(c-1-n_{1})\leq 1+(s-b)(c-2)$, as $n_{1}\geq 1$ and $s\geq b+2$. By Claim 2, the graph $\Gamma$ has the SPLS(b+1,s)(b+1,s) property, and hence we have $|\ell|\geq a_{1}+2-(s-1)b$. Hence $\beta+\alpha(r-1)-(s-1)b+1=a_{1}+2-(s-1)b\leq 1+(s-b)(c-2)$ and therefore $\beta\leq(s-b)(c-2+b)-\alpha(r-1)+b^{2}-b$. So, if $\beta>(s-b)(c-2+b)-\alpha(r-1)+b^{2}-b$, the sets $x,y_{1}]$ and $x,y_{2}]$ are equal and this shows Claim 5.
Claim 6 Assume that $x$ is a Delsarte vertex and that $y$ is a neighbour of $x$. Then $y$ is also a Delsarte vertex.
Proof of Claim 6. Let $C_{1},C_{2},\ldots,C_{r}$ be the lines through $x$ and without loss of generality, assume $y$ in $C_{1}$. The $C_{i}$’s are all Delsarte cliques by the fact that $x$ is a Delsarte vertex. Let $C_{1},D_{2},\ldots,D_{\sigma}$ be the lines through $y$. Then $\sigma\geq r$, and if equality holds, then $y$ is a Delsarte vertex. Let $\tilde{C}_{i}$ be the set of vertices in $C_{i}$ that are not neighbours of $y$, and $\tilde{D}_{i}$ be the set of vertices in $D_{i}$ that are not neighbours of $x$ for $i\geq 2$. Let $\Delta$ be the bipartite graph with vertex set $\{\tilde{C}_{2},\tilde{C}_{3},\ldots,\tilde{C}_{r},\tilde{D}_{2},\tilde{D}_{3},\ldots,\tilde{D}_{\sigma}\}$, and $\tilde{C}_{i}$ is adjacent to $\tilde{D}_{j}$ if and only if there are vertices $u\in\tilde{C}_{i}$ and $v\in\tilde{D}_{j}$ that are adjacent, for $i,j\geq 2$. We will obtain two bounds on the number of edges of $\Delta$ by double-counting. Let $v_{1},v_{2}$ be two distinct vertices of $\tilde{D}_{j}$ for some $j\geq 2$. By Claims 4 and 5, we have $x,v_{1}]=[x,v_{2}]$ both with cardinality $b+1$. As ${C}_{1}\in[x,v_{1}]$, we have that the valency $k_{j}$ of $\tilde{D}_{j}$ is at most $b$. Let $C_{i}\in[x,v_{1}]$ for some $i\geq 2$. Then, as $C_{i}$ is a Delsarte clique, there are $\alpha+1$ neighbours of $v_{1}$ inside $C_{i},$ by Lemma 10. Note that $|\Gamma(y)\cap C_{i}-\{x\}|=\alpha$, therefore $\Gamma(v_{1})\cap\tilde{C_{i}}\neq\emptyset$. This means that $k_{j}$ is equal to $b$ and the number $\varepsilon$ of edges of $\Delta$ satisfies $\varepsilon=b(\sigma-1).$ Let $u_{1},u_{2}$ be two distinct vertices of $\tilde{C}_{i}$ for some $i\geq 2$. Again by Claims 4 and 5, we have $y,u_{1}]=[y,u_{2}]$ both with cardinality $b+1$. As ${C}_{1}\in[y,u_{1}]$ we have that the valency $\kappa_{i}$ of $\tilde{C}_{i}$ is at most $b$. Therefore $\varepsilon\leq(r-1)b$ and hence $\sigma\leq r$. This means that $y$ is also a Delsarte vertex.
By Proposition 19 we are done as we have at least one Delsarte vertex and hence all the vertices must be Delsarte vertices. This shows that $\Gamma$ is geometric.
For geometric distance-regular graphs with classical parameters, we can provide a better bound for $\beta$ than the one in Claim 5 of Theorem 21, using the same discussion as before.
Theorem 22
Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$, having classical parameters $(D,b,\alpha,\beta)$, where $b\geq 2$ and $D\geq 3$. Let $r=[D]$. If $\beta>(r-\alpha-1)\alpha b+\alpha$, then for each pair of a vertex $x$ and a line $\ell$, with $d(x,\ell)=1$ and vertices $y_{1},y_{2}$ on $\ell$ with $d(x,y_{1})=d(x,y_{2})=2$, we have that the sets of lines $x,y_{1}]$ and $x,y_{2}]$ are equal.
Proof. Let $x,y$ be two vertices of $\Gamma$ at distance 2. We will first show $|[x,y]|=b+1$. If $\ell$ is a line in $[x,y]$, then $d(y,\ell)=1$ as $\ell$ is a Delsarte clique, and hence $|[x,y]|=\tau_{2}=b+1.$ Let $u,v$ be two vertices at distance 3 in $\Gamma$, and $m$ a line through $u$ and at distance 2 from $v$. Then the number of vertices $z$ at distance 2 from $v$ such that $z\in m$ is equal to $\phi_{2}=1+\alpha(b+1)$ and hence at most $b^{2}+b+1$.
Let $\ell$ be a line at distance 1 from $x$. Assume that there are vertices $y_{1},y_{2}$ at distance 2 from $x$ that are both on $\ell$ such that $x,y_{1}]\neq[x,y_{2}]$. Put, for $i=1,2$, $S_{i}:=\Gamma_{i}(x)\cap\ell$ and $n_{i}:=|S_{i}|$. By Lemma 10 we have $|\ell|=\beta+1$, $n_{1}=\phi_{1}=\alpha+1$ and $n_{2}=|\ell|-n_{1}=\beta-\alpha$, as $\Gamma$ is geometric. This means that there are $\alpha+1$ lines through $x$ that intersect $\ell$, as two distinct lines intersect in at most one point and $n_{1}=\alpha+1$. Those lines are in $T:=[x,y_{1}]\cap[x,y_{2}]$. Let $t:=|T|$. As $|[x,y_{1}]|=|[x,y_{2}]|=b+1$, and $x,y_{1}]\neq[x,y_{2}]$, we have $\alpha+1=n_{1}\leq t\leq b$. As there are exactly $r$ lines through $x$, there are exactly rtr-t lines through $x$ that are not in $T$. Let $u\sim x$ such that $d(y_{1},u)=3$. Then the line $\ell_{1}$ through both $u$ and $x$ does not lie in $x,y_{1}]$. Then $d(u,\ell)=2$ and hence $|\Gamma_{2}(u)\cap\ell|=\phi_{2}=1+\alpha(b+1)$, by Lemma 10. As $u\sim x$ and $d(u,y_{1})=3$, we find that every vertex in $S_{1}$ is at distance 2 from $u$. As a consequence $|\Gamma_{2}(u)\cap S_{2}|\leq|\Gamma_{2}(u)\cap\ell|-|S_{1}|=1+\alpha(b+1)-(1+\alpha)=\alpha b$. It follows that there are at most $\alpha b$ vertices $v$ in $S_{2}$ such that $\ell_{1}\in[x,v]$, as this implies $d(u,v)\leq 2$.
On the other hand, for $v\in S_{2}$, the set $[x,v]$ contains at least b+1tb+1-t lines through $x$ that are not in $T$. By counting pairs $(v,\ell_{2})$ with $v\in S_{2}$ and lines $\ell_{2}$ in $[x,v]$ that are not in $T$ we have $\beta-\alpha\leq(\beta-\alpha)(b+1-t)=n_{2}(b+1-t)\leq(r-t)\alpha b$, as $\ell_{2}$ contains a vertex $w$ that is in $\Gamma_{3}(y_{1})\cup\Gamma_{3}(y_{2})$ and $t\leq b$.
Therefore $\beta-\alpha\leq(r-t)\alpha b$. So, if $\beta>(r-\alpha-1)\alpha b+\alpha$, the sets $x,y_{1}]$ and $x,y_{2}]$ are equal. This shows the theorem.
In the next theorem we give a bound on $\beta$ to insure that a distance-regular graph with classical parameters is geometric.
Theorem 23
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$ such that $b\geq 2$ and $D\geq 3$. Let $r=[D]$. If
$\beta\geq\max\{\frac{2b+4}{2b+3}r(b+2)(\alpha b+b+\alpha),\frac{2b+4}{2b+3}r((b+1)(b^{2}+b+2)-3)\},$
then $\Gamma$ is geometric.
Proof. We only need to show the existence a positive integer $s$ such that the conditions in both Theorem 15 and Theorem 21 are satisfied. As we already have seen in the proof of Theorem 20, Equations (4) and (5), the conditions in Theorem 15 can be written as
$\beta>\frac{s+1}{2(s+1-r)}\left((\alpha b+\alpha+b)s-2\alpha(r-1)\right),$ (10)
and
$\beta>(\alpha b+b+\alpha)2s-\alpha r-(\alpha+1)b.$ (11)
Let $s_{0}:=\frac{2b+4}{2b+3}r$ be a rational number and $s:=\lfloor s_{0}\rfloor$. Then $\frac{s+1}{s+1-r}<\frac{s_{0}}{s_{0}-r}=2(b+2).$ So to show that Conditions (10) and (11) hold, it suffices to show that the following two conditions hold:
$\beta>(b+2)\left((\alpha b+b+\alpha)s_{0}-2\alpha(r-1)\right),$ (12)
and
$\beta$ $>2(\alpha b+b+\alpha)s_{0}-\alpha r-(\alpha+1)b.$ (13)
But these two conditions are satisfied as
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
This shows that $\Gamma$ satisfies the SPLS(s)(s) property. Now we need to consider the conditions in Theorem 21. These are
$\beta>(b+2)(b+1)(s-1)-(b(b+1)+r-1)\alpha,$ (14)
$\beta>((b+1)(b^{2}+b+1)-2)(s-b)-\alpha(b+1)r+\alpha(b+1)^{2},$ (15)
and
$\beta>((b+1)(b^{2}+b+1)-2+b)(s-b)-\alpha(r-1)+b^{2}-b.$ (16)
As
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}$
and $s_{0}=\frac{2b+4}{2b+3}r$, we find that
$\beta>(b+2)(b+1)(s_{0}-1)-(b(b+1)+r-1)\alpha,$ (17)
$\beta>((b+1)(b^{2}+b+1)-2)(s_{0}-b)-\alpha(b+1)r+\alpha(b+1)^{2},$ (18)
and
$\beta>((b+1)(b^{2}+b+1)-2+b)(s_{0}-b)-\alpha(r-1)+b^{2}-b$ (19)
are satisfied and hence the conditions of Theorem 21 are satisfied. This shows that $\Gamma$ is geometric and this finishes the proof.
As a corollary we have the following consequence for distance-regular graph with classical parameters $(D,b,0,\beta)$.
Corollary 24
Let $\Gamma$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$, such that $b\geq 2$, $\alpha=0$ and $D\geq 3$. Let $r=[D]$. Then $\beta<2\frac{b+2}{2b+3}(b^{3}+2b^{2}+3b-1)r$ holds.
Proof. For $\beta=1$ the corollary is clearly true, so we may assume $\beta>1$. If $\beta\geq 2\frac{b+2}{2b+3}(b^{3}+2b^{2}+3b-1)r$, then $\Gamma$ is geometric by Theorem 23. By Proposition 12, we find that $\Gamma$ is a dual polar graph or a Hamming graph. In the first case we have $\beta\leq b^{2}$, by [3, Table 6.1]. In the last case we have $b=1$. So the theorem follows.
6 The ELS property and strongly regular subgraphs
Let $\Gamma$ be a geometric distance-regular graph with respect to a set $\mathcal{C}$ of Delsarte cliques. We say that $\Gamma$ has the equal line set (ELS) property if the following condition holds: In the partial linear space $X=(V(\Gamma),\mathcal{C},\in)$, for each pair $x\in V(\Gamma)$ and $\ell\in\mathcal{C}$ such that $d(x,\ell)=1$, and for every pair of distinct vertices $y_{1}$ and $y_{2}$ on $\ell$ at distance $2$ from $x$, the sets of lines $x,y_{1}]$ and $x,y_{2}]$ are equal.
Lemma 25
Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$, having diameter $D\geq 3$. Assume that $\Gamma$ satisfies the ELS property. Let $x,y$ be two vertices of $\Gamma$ at distance 2 in $\Gamma$. Let $x,y]=\{\ell_{1},\ell_{2},\ldots,\ell_{\tau_{2}}\}$ and $y,x]=\{m_{1},m_{2},\ldots,m_{\tau_{2}}\}$. Let $z\in\Gamma(y)\cap\Gamma_{2}(x)$. Then $C_{2}(x,z)\subseteq\ell_{1}\cup\ell_{2}\cup\cdots\cup\ell_{\tau_{2}}$ if and only if there exists $i$ with $1\leq i\leq\tau_{2}$ such that $z\in m_{i}$.
Proof. If $z\in m_{i}$, then by the ELS property $x,y]=[x,z]=\{\ell_{1},\ell_{2},\ldots,\ell_{\tau_{2}}\}$ and we are done. Now assume that $C_{2}(x,z)\subseteq\ell_{1}\cup\ell_{2}\cup\cdots\cup\ell_{\tau_{2}}$. If $z$ is not on $m_{1}$, then there exists a vertex $u$ on $m_{1}$ adjacent to $x$, but not adjacent to $z$ as $z$ has exactly $\phi_{1}$ neighbours on $m_{1}$ including $y$, and $x$ has has exactly $\phi_{1}$ neighbours on $m_{1}$. Hence there exists a vertex $w\in C_{2}(x,z)$ not adjacent to $y$ as both $y$ and $z$ have $c_{2}$ common neighbours with $x$. Now by the ELS property we have $[y,x]=[y,w]$, and the line $L$ through $y$ and $z$ lies at distance 1 from $w$. So $L\in\{m_{1},m_{2},\ldots,m_{\tau_{2}}\}$. This shows that $z$ lies on $m_{i}$ for some $i=1,2,\ldots,\tau_{2}$. This shows the lemma.
Theorem 26
Let $\Gamma$ be a geometric distance-regular graph with respect to $\mathcal{C}$, having valency $k$, smallest eigenvalue $\theta_{\min}$, diameter $D\geq 3$, and $\phi_{1}\geq 2$. Let $r=-\theta_{\min}$ and $\beta=\frac{k}{-\theta_{\min}}$. Assume that $\Gamma$ satisfies the ELS property. Then the following hold: Let $x,y$ be a pair of vertices at distance 2 in $\Gamma$. Assume that $x,y]=\{\ell_{1},\ell_{2},\ldots,\ell_{\tau_{2}}\}$. Let $\Sigma=\Sigma(x,y)$ be the subgraph induced on $\{x\}\cup\ell_{1}\cup\ell_{2}\cup\cdots\cup\ell_{\tau_{2}}\cup\{z\in\Gamma_{2}(x)\mid[x,y]=[x,z]\}$. Then the graph $\Sigma$ has the following properties:
  1. (1)
    $\Sigma$ has diameter 2;
  2. (2)
    $\Sigma$ is geodetically closed;
  3. (3)
    If $\Sigma$ contains two vertices of a Delsarte clique $C\in{\mathcal{C}}$, then $\Sigma$ contains all the vertices of $C$;
  4. (4)
    $\Sigma$ is a geometric strongly regular graph with parameters
    $\left(\frac{\beta(\beta-\phi_{1}+1)(\tau_{2}-1)}{\phi_{1}}+\beta\tau_{2}+1,\ \beta\tau_{2},\ \beta-1+(\phi_{1}-1)(\tau_{2}-1),\ \phi_{1}\tau_{2}\right);$
  5. (5)
    For every pair of vertices at distance 2, there exists a unique induced subgraph $\Delta$ of $\Gamma$ containing them such that $\Delta$ has diameter 2, is geodetically closed, and has the property that whenever $\Delta$ contains two vertices of a Delsarte clique in $\mathcal{C}$, it contains the entire Delsarte clique.
Proof. By Lemma 3, the number $\beta=\frac{k}{-\theta_{\min}}$ is a positive integer and each vertex lies on $r$ lines as $\Gamma$ is geometric.
Let $x,y$ be two vertices at distance 2 in $\Gamma$. Let $\ell_{1},\ell_{2},\ldots,\ell_{r}$ respectively $m_{1},m_{2},\ldots,m_{r}$ be the lines through $x$ respectively $y$. Assume $x,y]=\{\ell_{1},\ell_{2},\ldots,\ell_{\tau_{2}}\}$ and $y,x]=\{m_{1},m_{2},\ldots,m_{\tau_{2}}\}$. Let $\Sigma$ be the graph as defined in the theorem. Then we have $y\in V(\Sigma)$ and $\{x\}\cup\Sigma(x)=\ell_{1}\cup\ell_{2}\cup\cdots\cup\ell_{\tau_{2}}$. By Lemma 25 we have that the neighbourhood of $z\in\Sigma_{2}(x)$ in $\Sigma$, denoted by $\Sigma(z)$ is the set $(\bigcup_{\ell\in[z,x]}\ell)\setminus\{z\}$. This, in particular, means that the valency of $z$ in $\Sigma$, denoted by $k_{\Sigma}(z)$ satisfies $k_{\Sigma}(z)=\beta\tau_{2}$.
Claim 1 Let $u$ be a vertex of $\Sigma$, not adjacent to $y$. Then $d_{\Sigma}(u,y)=2$ and $[y,u]=[y,x]$.
Proof of Claim 1. If $u\sim x$, then it is true by the ELS property. So we may assume that $d(x,u)=2$. Assume that $u,x]=\{p_{1},p_{2},\ldots,p_{\tau_{2}}\}$. By Lemma 25, we have that the line $p_{i}$ lies completely inside $\Sigma$ for $i=1,2,\ldots,\tau_{2}$. If $y$ has a neighbour in each line $p_{i}$ for $i=1,2,\ldots,\tau_{2}$ that is also adjacent to $x$, then it is evident that Claim 1 holds. Therefore, we may assume that there exists an $i=1,2,\ldots,\tau_{2}$ such that all neighbours of $y$ in $p_{i}$ are not adjacent to $x$. Note that $x$ has $\phi_{1}\geq 2$ neighbours in $p_{i}$ and they lie on $\phi_{1}$ lines of $x,y]=\{\ell_{1},\ell_{2},\ldots,\ell_{\tau_{2}}\}$. Without loss of generality we may assume that there are vertices $w_{j}$ on $\ell_{j}$ for $j=1,2$ such that $\{w_{1},w_{2}\}\subseteq p_{i}\cap\Sigma(x)$. By the ELS property, we have $y,w_{1}]=[y,w_{2}]=[y,x]=\{m_{1},m_{2},\ldots,m_{\tau_{2}}\}$. Let $w_{1},y]=\{q_{1},q_{2},\ldots,q_{\tau_{2}}\}$. Applying Lemma 25 to the tuple $(y,w_{1},\{m_{t}\}_{t=1}^{\tau_{2}},\{q_{t}\}_{t=1}^{\tau_{2}})$ in the role of $(x,y,\{\ell_{t}\}_{t=1}^{\tau_{2}},\{m_{t}\}_{t=1}^{\tau_{2}})$ shows that $w_{2}$ is on a line $q_{h}$ for some $1\leq h\leq\tau_{2}$. Note that the three vertices $u,w_{1},w_{2}$ all lie on $p_{i}$ and that $w_{1},w_{2}$ both lie on $q_{h}$. As each pair of vertices lie on at most one line we find that $p_{i}=q_{h}$ and therefore $u\in q_{h}$. This means that $d_{\Sigma}(u,y)=2$. As $u$ and $w_{1}$ are both at distance 2 from $y$ and lie on the line $q_{h}$, we find $y,u]=[y,w_{1}]=[y,x]$, by the ELS property. This shows Claim 1.
Claim 2 Let $u\in\Sigma_{1}(x)$. Then there exists $w\in\Sigma_{2}(x)$ such that $d(w,u)=2$.
Proof of Claim 2. Let $y\in\Sigma_{2}(x)$. Assume $y\sim u$, otherwise we are done. Consider the line $\ell$ through $x$ and $u$. Then $y$ has another neighbour $z$ on $\ell$ as $\phi_{1}\geq 2$. By Lemma 6, we have that $\phi_{1}\leq\frac{\beta+1}{2}$, as $D\geq 3$. This means that the vertex $u$ has at most $\phi_{1}-1\leq\frac{\beta-1}{2}$ neighbours on the line $m$ through $y$ and $z$ that are at distance two from $x$. But there are at least $\beta+1-\phi_{1}\geq\frac{\beta+1}{2}$ vertices on $m$ at distance 2 from $x$. So there exists a vertex $w$ on $m$ that is at distance 2 from both $u$ and $x$. This shows the claim.
By the definition of $\Sigma(x,y)$ we have $\Sigma(x,y)=\Sigma(x,z)$ for all $z\in\Sigma_{2}(x)$.
By Claim 1 we have $\Sigma(x,z)=\Sigma(z,x)$ for all $z\in\Sigma_{2}(x).$ Let $u\in\Sigma_{1}(x)$. Then there exists a vertex $w\in\Sigma_{2}(x)$ at distance 2 from $u$ by Claim 2. Hence $\Sigma(x,y)=\Sigma(x,w)=\Sigma(w,x)=\Sigma(w,u)=\Sigma(u,w)$ by Claims 1 and 2.
It is now easily verified that $\Sigma$ is strongly regular and satisfies properties (1), (2), and (3).
For property (4), it is clear that $\Sigma$ has the given parameters. By direct calculation, the smallest eigenvalue of $\Sigma$ is $\tau_{2}$, and hence a clique $C$ in $\Sigma$ is a Delsarte clique if and only if $|C|=\beta+1$. Therefore, each Delsarte clique $C\in\mathcal{C}$ with $C\subseteq V(\Sigma)$ is also a Delsarte clique in $\Sigma$. Together with property (3), this shows that $\Sigma$ is geometric.
To show the last item, let $\Delta$ be an induced subgraph satisfying the conditions of Item (5) containing two vertices $x$ and $y$ at distance 2. Then it follows easily that $\Delta$ is a strongly regular graph with the parameters as in Item (4). This means that for any vertex $z$ of $\Delta$ at distance $2$ of $x$ satisfies $[x,z]=[x,y]$ and hence $\Delta=\Sigma(x,y)$.
This shows the theorem.
Remark 3
Note that the idea of finding the kind of subgraphs as in Theorem 26 goes back to Shult and Yanushka [20] and Brouwer and Wilbrink [5]. The proof of Theorem 26 is inspired by ideas from Gavrilyuk and Koolen [8].
7 Proof of Theorem 1
In this section we give the proof of Theorem 1.
Proof of Theorem 1. Terwilliger classified the distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ with $D\geq 3$ and $b=1$, and showed that they are the graphs in Items (1)-(4), see [3, Theorem 6.1.1].
Therefore we may assume that $b\geq 2$. If $\alpha=0$, then we are done by Corollary 24. So we assume $\alpha>0$.
We first consider the case that $\alpha\in\{b,b-1\}$. Assume that $\Gamma$ is not a Grassmann graph or a bilinear forms graph. In this case, by [19, Corollary 1.3], we have: if $\alpha=b-1$, then
$\beta$ $<\frac{1}{2b-1}(2b^{4}+2b^{3}+2b^{2}+b-1)r$
$<\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\};$
if $\alpha=b$, then
$\beta$ $<\frac{8}{3}(b^{2}+2b)r$
$<\max\{\frac{2b+4}{2b+3}r(b+2)(\alpha b+b+\alpha),\frac{2b+4}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
From now on, we may assume that
$\beta\geq\max\{\frac{2}{2b+3}r(b+2)^{2}(\alpha b+b+\alpha),2\frac{b+2}{2b+3}r((b+1)(b^{2}+b+2)-3)\}.$
By Theorem 23, the graph $\Gamma$ is geometric and hence it has Delsarte clique. This means that $1\leq\alpha\leq b-2$ is an integer, as we already have considered $\alpha\in\{0,b-1,b\}$.
Note that $\Gamma$ is geometric. Let $\mathcal{C}$ be a set of Delsarte cliques of $\Gamma$ such that $\Gamma$ is the point graph of the partial linear space $X=(V(\Gamma),{\mathcal{C}},\in)$.
By Theorem 22, the graph $\Gamma$ has the ELS property. By Theorem 26, there exists an induced subgraph $\Sigma$ of $\Gamma$ such that $\Sigma$ is a strongly regular graph with classical parameters $(b,\alpha,\beta)$. Clearly $\Sigma$ is not a completely multipartite as the valency $k_{\Sigma}=(b+1)\beta>\mu(\Sigma)=(b+1)(\alpha+1)$ as the diameter of $\Gamma$ is at least 3. As $1\leq\alpha\leq b-2$, we see that $\mu(\Sigma)=(b+1)(\alpha+1)<(b+1)b$. Hence, by Theorem 13, we have $\beta\leq\frac{1}{2}(b+1)b((\alpha+1)(b+1)+1)+\alpha$. This gives a contradiction. This shows the theorem.
Acknowledgments
J.H. Koolen is partially supported by the National Natural Science Foundation of China (No. 12471335), and the Anhui Initiative in Quantum Information Technologies (No. AHY150000). We thank Bart De Bruyn and Ferdinand Ihringer for fruitful discussions.